### v2

# Junhan Kim    Ho-young Kim    Eulrang Cho    Chungman Lee    Joonyoung Kim    Yongkweon Jeon

Abstract
Quantization is a promising solution for deploying large language models (LLMs) on resource-constrained devices. Early methods developed for smaller networks like ResNet rely on gradient-based optimization, which becomes impractical for hyper-scale LLMs with billions of parameters. While recently proposed backpropagation-free or transformation-based methods alleviate this issue, their performance remains limited by either a lack of inter-layer dependency consideration or the use of naïve nearest-rounding-based integer weight assignment to save the heavy computational cost of weight optimization. We thus introduce a novel backpropagation-free PTQ algorithm that optimizes integer weights by considering inter-layer dependencies. The key innovation is the development of attention-aware Hessian matrices that capture inter-layer interactions within the attention module. Extensive experiments demonstrate that our approach not only outperforms existing weight quantization methods but also shows good synergy with conventional methods to suppress activation outliers, leading to state-of-the-art weight-activation quantization performance.
Machine Learning, ICML
Samsung Research
{jun_one.kim, dragwon.jeon}@samsung.com
1 Introduction
The explosive growth in complexity (parameters) of large language models (LLMs) (Touvron et al., 2023a; Zhang et al., 2022) has resulted in a proportional increase in computational costs, which has prompted an urgent need for efficient model processing and compression strategies. Quantization has emerged as a pivotal solution and an essential step for deploying AI models on resource-constrained devices that primarily support fixed-point arithmetic.
Two main categories of quantization approaches have been proposed to preserve the performance of original models: quantization-aware training (QAT) and post-training quantization (PTQ). Although QAT can potentially outperform PTQ, its practicality diminishes when handling hyper-scale LLMs featuring billions of parameters. Consequently, recent quantization efforts have been directed toward PTQ.
Although classic PTQ methods have successfully quantized small models such as ResNet (Nagel et al., 2020; Li et al., 2021), they rely on time-consuming gradient-based optimization, so their efficacy decreases when the complexity of LLMs increases. Accordingly, recent efforts have shifted to develop (a) backpropagation-free methods that optimize integer weights based on Hessian (Frantar et al., 2023) or (b) transformation-based methods that convert a model to be robust to quantization by applying smoothing, rotation, and permutation (Shao et al., 2023; Ashkboos et al., 2024; Liu et al., 2024; Lin et al., 2024a). However, their weight quantization performance is limited as they ignore inter-layer dependencies or use naïve nearest rounding for weight quantization.
In this paper, we propose a novel backpropagation-free PTQ algorithm that considers inter-layer dependencies in optimizing integer weights. Our contributions are as follows:
  • We propose a novel PTQ algorithm called BoA111 Backpropagation-free optimization for Attention-aware PTQ. The key contribution is to exploit the attention reconstruction error, not the layer-wise reconstruction error, in approximating the Hessian to consider inter-layer dependencies within the attention module (Section 3.2). To our knowledge, BoA is the first method to optimize integer weights by considering inter-layer dependencies without relying on gradient-based optimization.
  • While the proposed Hessian facilitates the consideration of inter-layer dependencies, it requires additional memory and computations. To mitigate the overhead, we propose several techniques, including Hessian relaxation, efficient computation of inverse Hessians, and head-wise simultaneous quantization (Section 3.3).
  • We evaluate BoA on publicly available LLMs. From extensive experiments, we show that BoA outperforms existing backpropagation-free weight quantization methods by a significant margin, particularly for low-bit precision (e.g., INT2) (Section 4.2). Furthermore, when combined with existing methods to suppress outliers (Ashkboos et al., 2024; Liu et al., 2024), BoA achieves the state-of-the-art performance for both weight-only and weight-activation quantization (Sections 4.2 and 4.3).
2 Related Works
When calibration data are available, PTQ aims to minimize the increase in task loss incurred by quantization. Assuming the convergence of a network and the independence between layers, the problem of quantizing weights to minimize task loss degradation can be formulated as the following layer-wise reconstruction problem (Nagel et al., 2020):
$\min_{\Delta\mathbf{W}^{(\ell)}}~{}$ $\left\|\left({\mathcal{Q}}(\mathbf{W}^{(\ell)})-\mathbf{W}^{(\ell)}\right)\mathbf{X}^{(\ell-1)}\right\|_{F}^{2},$ (1)
where $\mathbf{X}^{(\ell-1)}$ is the input to the $\ell$-th layer parameterized by $\mathbf{W}^{(\ell)}$ and ${\mathcal{Q}}$ is a quantization function. For a uniform quantization, if the nearest quantization bin is assigned to each weight, ${\mathcal{Q}}$ is defined as
${\mathcal{Q}}(x)$ $=s\left(\text{clamp}\left(\left\lfloor\frac{x}{s}\right\rceil+z,0,2^{n}-1\right)-z\right),$
[…]
Early works aimed to optimize the weight-rounding mechanism (Nagel et al., 2020; Hubara et al., 2021; Li et al., 2021). Instead of allocating the nearest integer, these studies attempted to assign integer weights that minimize the reconstruction error. One popular approach is AdaRound, which learns integer weights satisfying 1 via backpropagation (Nagel et al., 2020). This algorithm has been extended to Brecq where the block-wise reconstruction error has been used, instead of the layer-wise reconstruction error, to consider the inter-layer dependencies (Li et al., 2021). Although AdaRound and Brecq have successfully quantized small-sized models, they rely on time-consuming gradient-based optimizations, which renders their application to LLMs with billions of parameters challenging. Consequently, recent efforts have shifted towards the development of cost-effective quantization methods for LLMs.
These efforts can be classified into two orthogonal classes: (a) backpropagation-free Hessian-based integer weight optimization methods (e.g., GPTQ (Frantar et al., 2023)) and (b) transformation-based methods that convert a model to be more robust to quantization by applying smoothing (e.g., SmoothQuant (Xiao et al., 2023), OmniQuant (Shao et al., 2023), and AffineQuant (Ma et al., 2024)), rotation (e.g., QuaRot (Ashkboos et al., 2024) and SpinQuant (Liu et al., 2024)), or permutation (e.g., DuQuant (Lin et al., 2024a)). The proposed BoA belongs to the class (a) in that it optimizes integer weights using the Hessian without relying on backpropagation. Furthermore, similar to GPTQ, BoA can be combined with transformation-based methods such as QuaRot and SpinQuant (see Section 4). The primary difference over GPTQ lies in our optimization objective: while GPTQ assumes layer-wise independence and focuses on layer-wise reconstruction (which leads to performance degradation), our approach explicitly aims to preserve the attention output, enabling us to account for inter-layer dependencies within the attention module.
Other algorithms exploiting different strategies have also been proposed. For example, SpQR (Dettmers et al., 2023), SqueezeLLM (Kim et al., 2023), and OAC (Edalati et al., 2024) proposed a mixed-precision approach that assigns a large bit-width to quantization-sensitive weights or retains them in full-precision. Compared to the standard uniform quantization, these algorithms require additional processing and memory costs in the inference and need special hardware and dedicated kernels without which accelerating the inference may not be easy. Furthermore, unlike server-grade GPUs, on-device NPUs (e.g. Qualcomm Hexagon) lack support for the mixed precision format, and customizing kernels for desired functionalities is very challenging. Thus, we exclude these algorithms in our comparison and focus on the more universally supported uniform quantization format.
3 Method
3.1 Overview of Proposed BoA
The proposed BoA quantizes weights by repeating quantization and weight-update steps; once BoA quantizes one weight, it updates the remaining (not-yet-quantized) weights to compensate for the task loss degradation caused by the quantization. The update formula to compensate for the quantization of the $q$-th weight $w_{q}$ is formulated as (Frantar et al., 2023)
$\bm{\delta}{\bm{w}}=\frac{{\mathcal{Q}}(w_{q})-w_{q}}{[\mathbf{U}]_{q,q}}[\mathbf{U}]_{q,:}\text{ where }\mathbf{U}$ $=\operatornamewithlimits{Chol}(\mathbf{H}^{-1})^{T},$ (2)
where $\mathbf{H}$ is the Hessian and $\operatornamewithlimits{Chol}(\cdot)$ denotes a Cholesky decomposition (i.e., $\mathbf{U}$ is an upper triangular matrix satisfying $\mathbf{H}^{-1}=\mathbf{U}^{T}\mathbf{U}$).
The key difference over GPTQ lies in the approximation of the Hessian $\mathbf{H}$. In GPTQ, the layer-wise reconstruction error $\|\Delta\mathbf{W}^{(\ell)}\mathbf{X}^{(\ell-1)}\|_{F}^{2}$ in 1 has been used (i.e., the layer-wise independence has been assumed) to approximate $\mathbf{H}$ which yields the following Hessian equation [^2] The second-order derivative of $\|\mathbf{M}_{1}\Delta\mathbf{W}\mathbf{M}_{2}\|_{F}^{2}$ with respect to $\Delta\mathbf{w}$ is $2\mathbf{M}_{2}\mathbf{M}_{2}^{T}\otimes\mathbf{M}_{1}^{T}\mathbf{M}_{1}$ (see Appendix A for the proof). :
$\mathbf{H}^{({\bm{w}}^{(\ell)})}$ $\approx 2\mathbf{X}^{(\ell-1)}\mathbf{X}^{(\ell-1)^{T}}\otimes\mathbf{I},$ (3)
where $\mathbf{H}^{({\bm{w}}^{(\ell)})}$ is the Hessian for the $\ell$-th layer, $\otimes$ denotes the Kronecker product operation and $\mathbf{I}$ is the identity matrix. The approximated Hessian $\mathbf{H}^{({\bm{w}}^{(\ell)})}$ in GPTQ relies solely on the input, which means that GPTQ cannot consider the influence of other layers. In other words, GPTQ neglects inter-layer dependencies within the attention module, a crucial aspect of Transformers, which results in limited performance (Jeon et al., 2023). To overcome this, we develop Hessians that incorporate inter-layer dependencies and then use them instead of the conventional Hessian in 3.
In Table 1, we summarize the proposed Hessians; the detailed derivation is provided in the following subsections. It can be observed that the proposed Hessians not only contain the term related to the input $\mathbf{X}$, but also involve the terms related to other layers (e.g., $\mathbf{K}_{h}^{T}\mathbf{K}_{h}$ for $\mathbf{W}_{Q}$).
Table 1: Approximated Hessians in GPTQ and the proposed BoA
Method Layer $\mathbf{H}=\mathbf{H}_{\text{col}}\otimes\mathbf{H}_{\text{row}}$
GPTQ $\mathbf{W}_{\{Q,K,V\}}$ $2\mathbf{X}\mathbf{X}^{T}\otimes\mathbf{I}$
BoA $\mathbf{W}_{Q,h}$ $2\mathbf{X}\mathbf{X}^{T}\otimes\mathbf{K}_{h}^{T}\mathbf{K}_{h}$
$\mathbf{W}_{K,h}$ $2\mathbf{X}\mathbf{X}^{T}\otimes\mathbf{Q}_{h}^{T}\mathbf{Q}_{h}$
$\mathbf{W}_{V,h}$ $2\mathbf{X}\mathbf{A}_{h}^{T}\mathbf{A}_{h}\mathbf{X}^{T}\otimes\mathbf{W}_{\text{out},h}^{T}\mathbf{W}_{\text{out},h}$
$\mathbf{W}_{\text{out},h}$ $2\mathbf{X}_{\text{out},h}\mathbf{X}_{\text{out},h}^{T}\otimes\mathbf{I}$
Algorithm 1 BoA
0:  weights $\mathbf{W}\in\mathbb{R}^{d_{\text{row}}\times d_{\text{col}}}$ and inputs $\mathbf{X}$ of the Transformer layer
1:  Initialize quantized output: $\mathbf{Q}\leftarrow\mathbf{0}_{H\times d_{\text{row}}/H\times d_{\text{col}}}$
2:  Initialize (row-wise) quantization errors: $\mathbf{E}\leftarrow\mathbf{0}_{H\times d_{\text{col}}}$
3:  Compute attention-aware Hessians: $\mathbf{H}_{h}=\mathbf{H}_{\text{col},h}\otimes\mathbf{H}_{\text{row},h}$                           $\triangleright$ see Table 1
4:  Set step size (scale) $\mathbf{S}$: $\min_{\mathbf{S}}\hskip 0.56905pt\text{tr}\left(\Delta\mathbf{W}\mathbf{H}_{\text{col},h}\Delta\mathbf{W}^{T}\right)$
5:  Compute inverse Hessians $\mathbf{H}_{\text{col},h}^{-1}$ and $\mathbf{H}_{\text{row},h}^{-1}$
6:  Compute $\mathbf{U}_{\text{col},h}=\operatornamewithlimits{Chol}(\mathbf{H}_{\text{col},h}^{-1})^{T}$ and $\mathbf{U}_{\text{row},h}=\operatornamewithlimits{Chol}(\mathbf{H}_{\text{row},h}^{-1})^{T}$
7:  for $j=0,\ldots,d_{\text{row}}/H-1$ do
8:     Construct $\mathbf{W}^{(j)}\in\mathbb{R}^{H\times d_{\text{col}}}$ by stacking the $j$-th rows $[\mathbf{W}_{h}]_{j,:}$
9:     Quantize $\mathbf{W}^{(j)}$: $(\mathbf{Q}_{:,j,:},\mathbf{E})\leftarrow\text{GPTQ}(\mathbf{W}^{(j)},\mathbf{U}_{\text{col},h},\mathbf{S})$                            $\triangleright$ see Appendix D
10:     Update remaining rows: $[\mathbf{W}_{h}]_{j:,:}\leftarrow[\mathbf{W}_{h}]_{j:,:}-\frac{[\mathbf{U}_{\text{row},h}^{T}]_{j:,j}\cdot\mathbf{E}_{h,:}\cdot\mathbf{U}_{\text{col},h}}{[\mathbf{U}_{\text{row},h}]_{j,j}}$    $\triangleright$ see Proposition 3.1
11:  end for
11:  quantized weights $\mathbf{Q}$
3.2 Proposed Attention-aware Hessian
To consider the inter-layer dependencies within the attention module, we exploit the attention reconstruction error rather than the layer-wise reconstruction error when approximating the Hessian. For an input sequence $\mathbf{X}\in\mathbb{R}^{d\times L}$, the output of the multi-head attention (MHA) is expressed as
$\operatornamewithlimits{MHA}(\mathbf{X})=\sum_{h=1}^{H}\mathbf{W}_{\text{out},h}(\mathbf{A}_{h}\mathbf{V}_{h})^{T},~{}\mathbf{A}_{h}=\sigma\left(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{T}}{\sqrt{d_{h}}}\right),$
where $\mathbf{Q}_{h},\mathbf{K}_{h},\mathbf{V}_{h}\in\mathbb{R}^{L\times d_{h}}$ are query, key, value for the $h$-th attention head, $d_{h}$ is the head dimension, $\sigma$ is the row-wise softmax function, and $H$ is the total number of heads.
Hessian for query
When quantizing the query projection weights $\mathbf{W}_{Q,h}$, $\mathbf{W}_{\text{out},h}$ and $\mathbf{V}_{h}$ remain unchanged, but the attention weights $\mathbf{A}_{h}$ change. Using the first-order Taylor polynomial, the perturbation in $\mathbf{A}_{h}$ can be approximated as
$\Delta\mathbf{A}_{h}$ $=\sigma\left(\frac{(\mathbf{Q}_{h}+\Delta\mathbf{Q}_{h})\mathbf{K}_{h}^{T}}{\sqrt{d_{h}}}\right)-\sigma\left(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{T}}{\sqrt{d_{h}}}\right)$
$\approx\frac{\Delta\mathbf{Q}_{h}\mathbf{K}_{h}^{T}}{\sqrt{d_{h}}}\mathbf{J}_{\sigma}^{T}=\frac{\mathbf{X}^{T}\Delta\mathbf{W}_{Q,h}^{T}\mathbf{K}_{h}^{T}\mathbf{J}_{\sigma}^{T}}{\sqrt{d_{h}}},$ (4)
where $\mathbf{J}_{\sigma}$ is the Jacobian matrix of the softmax function $\sigma$. Thus, the attention reconstruction error is expressed as
$\|\Delta\operatornamewithlimits{MHA}(\mathbf{X})\|_{F}^{2}$ $=\|\mathbf{W}_{\text{out},h}(\Delta\mathbf{A}_{h}\mathbf{V}_{h})^{T}\|_{F}^{2}$
$=\left\|\frac{\mathbf{W}_{\text{out},h}\mathbf{V}_{h}^{T}\mathbf{J}_{\sigma}\mathbf{K}_{h}}{\sqrt{d_{h}}}\Delta\mathbf{W}_{Q,h}\mathbf{X}\right\|_{F}^{2},$ (5)
which yields the following Hessian for $\mathbf{W}_{Q,h}$ (see Footnote 2):
$\mathbf{H}^{(\mathbf{w}_{Q,h})}$ $=2\mathbf{X}\mathbf{X}^{T}\otimes\frac{\mathbf{K}_{h}^{T}\mathbf{J}_{\sigma}^{T}\mathbf{V}_{h}\mathbf{W}_{\text{out},h}^{T}\mathbf{W}_{\text{out},h}\mathbf{V}_{h}^{T}\mathbf{J}_{\sigma}\mathbf{K}_{h}}{d_{h}}.$ (6)
Hessian for key
As in the quantization of the query projection weights, the attention weight $\mathbf{A}_{h}$ changes when quantizing the key projection weights $\mathbf{W}_{K,h}$. By following the steps for 4, $\mathbf{A}_{h}$ can be approximated as
$\Delta\mathbf{A}_{h}$ $\approx\frac{\mathbf{Q}_{h}\Delta\mathbf{K}_{h}^{T}}{\sqrt{d_{h}}}\mathbf{J}_{\sigma}^{T}=\frac{\mathbf{Q}_{h}\Delta\mathbf{W}_{K,h}\mathbf{X}\mathbf{J}_{\sigma}^{T}}{\sqrt{d_{h}}},$
and then the attention reconstruction error is expressed as
$\|\Delta\operatornamewithlimits{MHA}(\mathbf{X})\|_{F}^{2}=\left\|\frac{\mathbf{Q}_{h}}{\sqrt{d_{h}}}\Delta\mathbf{W}_{K,h}\mathbf{X}\mathbf{J}_{\sigma}^{T}\mathbf{V}_{h}\mathbf{W}_{\text{out},h}^{T}\right\|_{F}^{2}.$
Thus, we obtain the following Hessian for $\mathbf{W}_{K,h}$:
$\mathbf{H}^{(\mathbf{w}_{K,h})}$ $=2\mathbf{X}\mathbf{J}_{\sigma}^{T}\mathbf{V}_{h}\mathbf{W}_{\text{out},h}^{T}\mathbf{W}_{\text{out},h}\mathbf{V}_{h}^{T}\mathbf{J}_{\sigma}\mathbf{X}^{T}\otimes\frac{\mathbf{Q}_{h}^{T}\mathbf{Q}_{h}}{d_{h}}.$ (7)
Hessian for value
When quantizing the weights $\mathbf{W}_{V,h}$ of the value projection, only $\mathbf{V}_{h}$ changes. Thus, we have
$\left\|\Delta\operatornamewithlimits{MHA}(\mathbf{X})\right\|_{F}^{2}$ $=\|\mathbf{W}_{\text{out},h}(\mathbf{A}_{h}\Delta\mathbf{V}_{h})^{T}\|_{F}^{2}$
$=\|\mathbf{W}_{\text{out},h}\Delta\mathbf{W}_{V,h}\mathbf{X}\mathbf{A}_{h}^{T}\|_{F}^{2},$
which yields the following Hessian for $\mathbf{W}_{V,h}$:
$\mathbf{H}^{(\mathbf{w}_{V,h})}$ $=2\mathbf{X}\mathbf{A}_{h}^{T}\mathbf{A}_{h}\mathbf{X}^{T}\otimes\mathbf{W}_{\text{out},h}^{T}\mathbf{W}_{\text{out},h}.$ (8)
Hessian for out
When the out-projection weights $\mathbf{W}_{\text{out},h}$ are quantized, the attention reconstruction error is
$\left\|\Delta\operatornamewithlimits{MHA}(\mathbf{X})\right\|_{F}^{2}$ $=\|\Delta\mathbf{W}_{\text{out},h}(\mathbf{A}_{h}\mathbf{V}_{h})^{T}\|_{F}^{2}.$
Thus, the corresponding Hessian is
$\mathbf{H}^{(\mathbf{w}_{out,h})}$ $=2\mathbf{V}_{h}^{T}\mathbf{A}_{h}^{T}\mathbf{A}_{h}\mathbf{V}_{h}\otimes\mathbf{I}=2\mathbf{X}_{\text{out},h}\mathbf{X}_{\text{out},h}^{T}\otimes\mathbf{I},$ (9)
where $\mathbf{X}_{\text{out},h}=(\mathbf{A}_{h}\mathbf{V}_{h})^{T}$.
3.3 Efficient Implementation of BoA
While inter-layer dependencies within the attention module can be considered by exploiting the proposed Hessians, they are significantly more complex than the conventional Hessian in 3, which may incur high computational costs. For example, computing the proposed Hessians in 6 and 7 would be more expensive than computing the conventional one in 3. In this subsection, we present techniques to mitigate the computational overheads incurred by the proposed attention-aware Hessians.
Hessian relaxation
The largest overhead related to the computation of the proposed Hessians is the Jacobian matrix $\mathbf{J}_{\sigma}$ in 6 and 7. For an input sequence of length $L$, the shape of $\mathbf{J}_{\sigma}$ is $H\times L\times L\times L$, which requires a large amount of memory and high computational cost (e.g., more than 400 GB even for the OPT-125M model when $L=2048$).
To mitigate such overhead, we establish a relaxed Hessian that does not require computing $\mathbf{J}_{\sigma}$. To this end, we build the following upper bound for the attention reconstruction error in 5, which will be used as its surrogate:
$\|\Delta\operatornamewithlimits{MHA}(\mathbf{X})\|_{F}^{2}$ $\leq\left\|\frac{\mathbf{W}_{\text{out},h}\mathbf{V}_{h}^{T}\mathbf{J}_{\sigma}}{\sqrt{d_{h}}}\right\|_{F}^{2}\cdot\left\|\mathbf{K}_{h}\Delta\mathbf{W}_{Q,h}\mathbf{X}\right\|_{F}^{2}.$
Noting that the constant term $\|\mathbf{W}_{\text{out},h}\mathbf{V}_{h}^{T}\mathbf{J}_{\sigma}\|_{F}^{2}$ does not affect quantization, [^3] The update $\bm{\delta}\mathbf{w}$ in 2 is not affected by the constant multiple of $\mathbf{H}$ because $[c\mathbf{U}]_{q,:}/[c\mathbf{U}]_{q,q}=[\mathbf{U}]_{q,:}/[\mathbf{U}]_{q,q}$ for any constant $c$. we use the term $\|\mathbf{K}_{h}\Delta\mathbf{W}_{Q,h}\mathbf{X}\|_{F}^{2}$ as a surrogate of the attention reconstruction error when deriving the Hessian for $\mathbf{W}_{Q,h}$, which yields the following Hessian:
$\mathbf{H}^{(\mathbf{w}_{Q,h})}$ $=2\mathbf{X}\mathbf{X}^{T}\otimes\mathbf{K}_{h}^{T}\mathbf{K}_{h}.$ (10)
Similarly, we can establish a relaxed Hessian for $\mathbf{W}_{K,h}$ as
$\mathbf{H}^{(\mathbf{w}_{K,h})}$ $=2\mathbf{X}\mathbf{X}^{T}\otimes\mathbf{Q}_{h}^{T}\mathbf{Q}_{h}.$ (11)
We summarize the relaxed Hessians in Table 1.
Efficient computation of inverse Hessians
After computing Hessians, their inverse matrices need to be computed to update the remaining weights after each quantization (see 2). Owing to the size of the proposed Hessians being $dd_{h}\times dd_{h}$, the complexity of the computation of the inverse Hessian would be ${\mathcal{O}}(d^{3}d_{h}^{3})$ in our approach. This is considerably more expensive than the complexity ${\mathcal{O}}(d^{3})$ of GPTQ, where the inverse of only $\mathbf{X}\mathbf{X}^{T}\in\mathbb{R}^{d\times d}$ in 3 is needed (Frantar et al., 2023).
For the efficient inverse computation, we exploit the useful properties of the Kronecker product (see 13a-13c in Appendix A). Specifically, let $\mathbf{H}=\mathbf{H}_{\text{col}}\otimes\mathbf{H}_{\text{row}}$ where $\mathbf{H}_{\text{col}}\in\mathbb{R}^{d\times d}$ and $\mathbf{H}_{\text{row}}\in\mathbb{R}^{d_{h}\times d_{h}}$, then we obtain
$\mathbf{H}^{-1}$ $=(\mathbf{H}_{\text{col}}\otimes\mathbf{H}_{\text{row}})^{-1}=\mathbf{H}_{\text{col}}^{-1}\otimes\mathbf{H}_{\text{row}}^{-1}.$
This implies that the inverse Hessian $\mathbf{H}^{-1}$ can be obtained by computing $\mathbf{H}_{\text{col}}^{-1}$ and $\mathbf{H}_{\text{row}}^{-1}$ (line 5 in Algorithm 1) whose complexity is ${\mathcal{O}}(d^{3})+{\mathcal{O}}(d_{h}^{3})$ ($={\mathcal{O}}(d^{3})$), not ${\mathcal{O}}(d^{3}d_{h}^{3})$. Similarly, we can efficiently perform the Cholesky decomposition (i.e., $\operatornamewithlimits{Chol}(\mathbf{H}^{-1})$ in 2) with the same order of complexity as in GPTQ. Specifically, if $\mathbf{L}_{1}=\operatornamewithlimits{Chol}(\mathbf{H}_{\text{col}}^{-1})$ and $\mathbf{L}_{2}=\operatornamewithlimits{Chol}(\mathbf{H}_{\text{row}}^{-1})$, $\mathbf{H}^{-1}$ can be expressed as
$\mathbf{H}^{-1}$ $=\mathbf{L}_{1}\mathbf{L}_{1}^{T}\otimes\mathbf{L}_{2}\mathbf{L}_{2}^{T}=(\mathbf{L}_{1}\otimes\mathbf{L}_{2})(\mathbf{L}_{1}\otimes\mathbf{L}_{2})^{T}.$
Subsequently, noting that the Kronecker product of lower triangular matrices is also lower triangular, we obtain
$\operatornamewithlimits{Chol}(\mathbf{H}^{-1})$ $=\mathbf{L}_{1}\otimes\mathbf{L}_{2}=\operatornamewithlimits{Chol}(\mathbf{H}_{\text{col}}^{-1})\otimes\operatornamewithlimits{Chol}(\mathbf{H}_{\text{row}}^{-1}).$
Thus, we can obtain $\operatornamewithlimits{Chol}(\mathbf{H}^{-1})$ by computing $\operatornamewithlimits{Chol}(\mathbf{H}_{\text{col}}^{-1})$ and $\operatornamewithlimits{Chol}(\mathbf{H}_{\text{row}}^{-1})$ (line 6 in Algorithm 1). Consequently, the computational complexity of the Choleksy decomposition would be ${\mathcal{O}}(d^{3})$, not ${\mathcal{O}}(d^{3}d_{h}^{3})$.
Simultaneous quantization of different heads
Because the proposed Hessians model the dependency between different rows (e.g., $\mathbf{H}_{\text{row}}=\mathbf{K}_{h}^{T}\mathbf{K}_{h}$ for $\mathbf{W}_{Q,h}$; see 10), we can compensate for the quantization error of a certain row by updating other rows. Specifically, the row-update formula is formulated as in the following proposition for given Hessian $\mathbf{H}=\mathbf{H}_{\text{col}}\otimes\mathbf{H}_{\text{row}}$.
Proposition 3.1.
Let $\mathbf{W}_{h}$ be a $d_{h}\times d$ matrix whose Hessian is $\mathbf{H}_{h}=\mathbf{H}_{\text{col},h}\otimes\mathbf{H}_{\text{row},h}$. Suppose $\mathbf{W}_{h}$ is quantized via 2. If the $j$-th row of $\mathbf{W}_{h}$ has been quantized, then the update formula to compensate for the quantization is expressed as
$[\delta\mathbf{W}_{h}]_{j:,:}=-\frac{[\mathbf{U}_{\text{row},h}^{T}]_{j:,j}\cdot\mathbf{e}\cdot\mathbf{U}_{\text{col},h}}{[\mathbf{U}_{\text{row},h}]_{j,j}},$ (12)
where $\mathbf{U}_{\text{col},h}=\operatornamewithlimits{Chol}(\mathbf{H}_{\text{col},h}^{-1})^{T}$, $\mathbf{U}_{\text{row},h}=\operatornamewithlimits{Chol}(\mathbf{H}_{\text{row},h}^{-1})^{T}$, and $\mathbf{e}\in\mathbb{R}^{1\times d}$ is the quantization error of the $j$-th row defined as $e_{i}=([\mathbf{W}_{h}]_{j,i}-\mathcal{Q}([\mathbf{W}_{h}]_{j,i}))/[\mathbf{U}_{\text{col},h}]_{i,i}$.
Proof.
See Appendix B. ∎
Figure 1: Illustration of BoA for the query projection $\mathbf{W}_{Q}$.
Proposition 3.1 implies that the conventional GPTQ cannot compensate quantization error of certain row by updating other rows because $\mathbf{H}_{\text{row}}=\mathbf{I}$ (see 3) and thus $\mathbf{U}_{\text{row}}=\mathbf{I}$ and $[\delta\mathbf{W}]_{i,:}=\mathbf{0}$ for all $i\neq j$.
While the proposed BoA can compensate for the quantization error of each row, the rows must be quantized sequentially (not simultaneously). For example, the second row can be quantized after being updated to compensate for the quantization error of the first row. To accelerate the quantization process, we assume independence between different attention heads (see Figure 1(a)), under which rows related to different heads are independent and can thus be quantized together. For a better understanding, we consider the query projection $\mathbf{W}_{Q}$ as an example (see Figure 1(b)). In the quantization step, we stack the $j$-th rows $[\mathbf{W}_{Q,h}]_{j,:}$ of all different heads, constructing the sub-weight matrix $\mathbf{W}_{Q}^{(j)}\in\mathbb{R}^{H\times d}$ (line 8 in Algorithm 1). Because the rows of $\mathbf{W}_{Q}^{(j)}$ are mutually independent, all the rows of $\mathbf{W}_{Q}^{(j)}$ can be quantized simultaneously as in GPTQ (line 9 in Algorithm 1). Following the quantization of $j$-th rows, we compensate for the quantization error by updating the remaining rows. In this update step, we use the refined weight-update formula in 12 (line 10 in Algorithm 1).
We measure quantization processing times of BoA with and without simultaneous quantization to evaluate how much the head-wise simultaneous quantization can accelerate the quantization process. From Table 2, we observe that BoA requires a significantly long processing time without the simultaneous quantization (more than one day even for the 7B model). This is because all rows need to be quantized sequentially (e.g., 4096 rows are quantized sequentially for LLaMA-7b) and thus the massive compute capabilities of modern GPUs cannot be utilized properly. As evident, we can significantly reduce the processing time by applying the head-wise simultaneous quantization (more than 40 times reduction on the 30B model).
Table 2: Processing time (hour) of BoA with and without simultaneous quantization of different heads
Simultaneous Quantization 7B 13B 30B
X 27.75 51.66 135.4
O 0.961 1.553 3.295
Table 3: Weight-only quantization performance on LLaMA2 and LLaMA3 models without transformation
Model Precision Method Wiki2 PPL ($\downarrow$) Zero-shot Accuracy ($\uparrow$)
Arc-c Arc-e BQ HS LAMB OBQA PIQA WG Average
LLaMA2-7B FP16 Baseline 5.473 45.90 74.66 77.92 75.94 70.86 44.00 78.89 68.90 67.13
INT2 RTN 7.8e3 26.45 26.18 39.30 25.99 0.00 24.20 49.40 49.96 30.19
GPTQ 30.85 26.96 44.15 56.30 42.52 23.95 28.20 63.22 54.70 42.50
BoA 12.76 28.33 47.73 64.71 45.46 31.33 29.60 64.85 54.46 45.81
INT3 RTN 342.4 22.53 35.06 44.13 28.87 1.63 26.40 58.11 48.70 33.18
GPTQ 6.719 36.95 59.72 65.54 66.68 58.21 39.20 74.65 66.06 58.38
BoA 6.007 40.27 69.28 74.22 72.00 66.60 41.20 78.35 67.64 63.70
LLaMA2-13B FP16 Baseline 4.885 49.06 77.65 80.49 79.38 73.41 45.80 80.69 72.22 69.84
INT2 RTN 5.7e3 27.47 26.89 37.95 25.97 0.00 25.20 49.08 48.38 30.12
GPTQ 35.08 21.76 35.31 61.96 35.22 19.41 28.80 57.29 52.88 39.08
BoA 18.33 29.78 46.30 62.23 49.30 29.33 27.60 63.33 52.64 45.06
INT3 RTN 227.2 23.98 28.75 49.24 28.51 4.46 24.40 53.26 50.75 32.92
GPTQ 9.790 34.81 62.12 67.06 55.57 47.61 37.00 73.18 61.33 54.84
BoA 5.833 43.52 69.28 78.93 74.71 65.22 35.20 77.42 62.51 63.35
LLaMA3-8B FP16 Baseline 6.137 53.67 77.61 81.19 79.15 72.23 45.00 81.01 73.24 70.39
INT2 RTN 6.6e4 25.51 25.80 53.94 26.34 0.00 29.00 51.52 50.20 32.79
GPTQ 24.54 23.29 32.58 53.39 39.17 7.10 27.00 53.32 52.25 36.01
BoA 21.70 26.62 44.87 60.52 43.34 16.89 29.40 59.85 55.80 42.16
INT3 RTN 129.1 23.21 33.88 55.60 34.83 4.60 25.20 58.22 52.57 36.01
GPTQ 8.226 41.47 63.01 75.47 70.05 59.53 40.60 73.78 69.85 61.72
BoA 7.782 45.14 72.77 78.69 72.66 62.41 42.60 77.31 71.35 65.37
LLaMA3.2-1B FP16 Baseline 13.15 38.14 63.26 69.51 60.78 54.38 34.60 74.37 59.51 56.82
INT2 RTN 6.3e4 26.96 25.59 41.53 26.05 0.01 26.40 51.52 50.59 31.08
GPTQ 538.9 25.26 26.64 37.83 26.41 0.22 27.60 51.41 48.46 30.48
BoA 312.2 25.09 26.85 40.06 27.17 1.42 27.00 51.96 51.07 31.33
INT3 RTN 1.9e3 25.60 26.94 54.13 29.90 0.58 27.00 52.18 49.17 33.19
GPTQ 112.0 24.06 39.48 53.85 31.07 13.85 27.80 60.07 49.33 37.44
BoA 26.43 30.63 55.26 59.97 48.33 31.95 29.60 66.65 53.99 47.05
LLaMA3.2-3B FP16 Baseline 11.04 46.16 67.80 78.62 70.44 62.15 36.00 75.52 67.40 63.01
INT2 RTN 2.0e4 26.79 26.52 37.89 25.93 0.00 30.80 50.92 49.09 30.99
GPTQ 98.19 24.83 27.78 52.32 33.83 4.38 28.60 52.23 51.14 34.39
BoA 54.64 25.77 35.48 57.52 35.63 14.42 29.00 56.96 53.43 38.53
INT3 RTN 882.6 26.37 27.86 45.87 37.04 1.44 26.00 53.92 48.46 33.37
GPTQ 46.14 28.92 37.63 44.01 39.27 18.76 28.60 61.64 54.70 39.19
BoA 13.64 42.32 66.12 77.52 64.46 54.18 35.20 72.69 62.51 59.38
    Results for high bit-widths and results on LLaMA1 and OPT models are provided in Section C.1 due to the page limitation.
Table 4: Weight-only quantization performance on transformed LLaMA2 and LLaMA3 models
Model Precision Method Wiki2 PPL ($\downarrow$) Zero-shot Accuracy ($\uparrow$)
Arc-c Arc-e BQ HS LAMB OBQA PIQA WG Average
LLaMA2-7B FP Baseline 5.473 45.90 74.66 77.92 75.94 70.86 44.00 78.89 68.90 67.13
INT2 OmniQuant 21.85 25.17 36.91 61.80 38.38 5.80 28.40 57.40 49.49 37.92
AffineQuant 91.19 23.46 29.17 40.03 27.44 0.16 22.20 54.13 51.78 31.05
QuaRot-RTN 1.1e4 27.22 26.05 37.83 26.19 0.01 25.60 50.54 48.70 30.27
QuaRot-GPTQ 22.05 22.27 36.24 61.99 33.54 20.63 28.60 56.75 51.78 38.98
DuQuant 1.6e4 26.62 25.72 39.08 26.06 0.00 22.20 49.73 50.12 29.94
DuQuant + LWC 46.27 23.12 28.96 44.04 29.60 2.41 27.40 53.70 49.17 32.30
BoA 10.42 29.69 54.84 64.37 52.03 46.54 33.60 67.52 59.43 51.00
LLaMA2-13B FP Baseline 4.885 49.06 77.65 80.49 79.38 73.41 45.80 80.69 72.22 69.84
INT2 OmniQuant 12.92 26.71 46.00 63.64 51.19 16.93 31.40 64.64 52.64 44.14
AffineQuant 9.415 26.96 49.83 63.15 52.40 26.58 33.60 64.47 53.83 46.35
QuaRot-RTN 7.9e3 27.22 26.26 37.83 25.74 0.00 24.00 49.02 49.17 29.91
QuaRot-GPTQ 9.593 31.66 56.52 63.15 48.81 36.93 32.60 66.05 60.38 49.51
DuQuant 1.4e4 28.33 26.64 44.43 26.12 0.00 25.00 50.16 49.57 31.28
DuQuant + LWC 10.40 28.50 46.51 63.70 52.79 25.24 30.80 65.13 54.14 45.85
BoA 8.237 35.49 63.97 71.28 58.36 56.25 35.40 71.82 62.75 56.92
LLaMA3-8B FP Baseline 6.137 53.67 77.61 81.19 79.15 72.23 45.00 81.01 73.24 70.39
INT2 OmniQuant 955.8 22.78 28.24 37.83 26.18 0.00 27.20 52.83 49.01 30.51
AffineQuant 1.1e3 23.46 27.31 37.86 26.04 0.00 26.80 52.12 51.22 30.60
QuaRot-RTN 3.5e5 26.19 25.21 39.02 26.86 0.00 28.60 50.38 49.41 30.71
QuaRot-GPTQ 18.28 28.33 46.68 64.10 45.04 29.49 29.20 62.08 55.25 45.02
DuQuant 1.4e6 25.00 25.34 47.37 26.68 0.00 29.80 51.96 48.93 31.89
DuQuant + LWC 2.6e4 24.40 26.68 38.26 25.30 0.00 29.20 49.67 52.01 30.69
BoA 15.24 30.38 54.42 68.17 49.17 39.89 34.20 65.94 60.14 50.29
LLaMA3.2-1B FP Baseline 13.15 38.14 63.26 69.51 60.78 54.38 34.60 74.37 59.51 56.82
INT2 OmniQuant 302.3 22.01 31.69 37.95 27.64 0.14 25.60 54.95 50.12 31.26
AffineQuant 268.3 22.44 31.61 37.83 27.75 0.12 26.00 54.30 49.88 31.24
QuaRot-RTN 2.6e5 26.62 25.84 38.75 26.75 0.00 28.20 51.09 51.07 31.04
QuaRot-GPTQ 54.28 22.18 34.01 56.39 31.90 14.65 24.80 56.91 50.59 36.43
DuQuant 4.9e4 25.51 25.97 39.94 26.17 0.00 29.80 48.97 49.64 30.75
DuQuant + LWC 9.3e3 28.16 25.38 37.89 25.27 0.00 26.60 50.49 49.57 30.42
BoA 40.86 24.06 39.77 58.20 34.62 16.79 26.20 57.07 52.64 38.67
LLaMA3.2-3B FP Baseline 11.04 46.16 67.80 78.62 70.44 62.15 36.00 75.52 67.40 63.01
INT2 OmniQuant 273.4 23.38 31.82 56.42 29.35 0.28 27.80 56.15 51.07 34.53
AffineQuant 282.2 22.35 32.87 47.28 29.44 0.25 28.20 56.26 48.38 33.13
QuaRot-RTN 2.3e4 26.19 24.49 47.09 26.55 0.00 30.40 51.74 48.30 31.85
QuaRot-GPTQ 52.18 23.98 36.95 60.92 34.68 19.24 27.40 57.67 52.49 39.17
DuQuant 7.7e4 25.17 25.55 41.10 26.01 0.00 28.40 51.47 50.28 31.00
DuQuant + LWC 770.9 23.63 26.64 38.20 26.32 0.03 26.20 51.74 51.70 30.56
BoA 33.40 27.30 45.66 65.87 40.86 24.57 29.00 61.37 56.27 43.86
  • $\dagger$
    BoA has been applied after transforming the model via QuaRot.
    The learnable equivalent transformation (LET) option has been deactivated because this option does not support models exploiting grouped-query attention (GQA).
    Results for high bit-widths and results on LLaMA1 models are provided in Section C.2 due to the page limitation.
4 Experiments
In this section, we evaluate the weight-only quantization performance of the proposed BoA (Section 4.2) and synergy with transformation-based methods in terms of weight-only and weight-activation quantization (Sections 4.2 and 4.3).
4.1 Experimental Setup
We conduct experiments on OPT (Zhang et al., 2022), LLaMA (Touvron et al., 2023a)), LLaMA2 (Touvron et al., 2023b), and LLaMA3. As a calibration dataset, we use 128 random sequences of length 2048 from WikiText-2 (Merity et al., 2016), as in (Shao et al., 2023; Ma et al., 2024; Lin et al., 2024a; Ashkboos et al., 2024; Liu et al., 2024). We report the perplexity (PPL) score on the WikiText-2 test dataset and accuracy on eight zero-shot commonsense reasoning tasks (ARC-challenge (Arc-c) and ARC-easy (Arc-e) (Clark et al., 2018), BoolQ (BQ) (Clark et al., 2019), HellaSwag (HS) (Zellers et al., 2019), LAMBADA (LAMB) (Paperno et al., 2016), OpenbookQA (OBQA) (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), and WinoGrande (WG) (Sakaguchi et al., 2021)). All experiments were conducted using a single NVIDIA H100 GPU (80 GB).
4.2 Weight-Only Quantization Results
Performance of integer weight optimization
We first evaluate the weight-only quantization performance of the proposed BoA. In this experiment, to solely compare the performance of the integer weight optimization, we do not apply any transform to the models, such as smoothing (Xiao et al., 2023; Lin et al., 2024b; Jeon et al., 2023; Shao et al., 2023; Ma et al., 2024), rotation (Ashkboos et al., 2024; Liu et al., 2024), or permutation (Lin et al., 2024a).
In Table 3 and Tables 7-9 (see Section C.1), we summarize the PPL and zero-shot task performances of BoA. For comparison, we include the performance of the conventional GPTQ and the rounding-to-nearest (RTN) method which naively assigns the nearest quantization bin. We observe BoA and GPTQ exhibit reasonable PPL even for INT2 quantization where RTN collapses significantly (i.e., PPL is almost $10^{4}$). This is because BoA and GPTQ minimize the task loss degradation rather than the weight quantization error $\Delta\mathbf{W}$. As evident, BoA significantly surpasses GPTQ on all models in both PPL and zero-shot accuracy. For example, BoA achieves 10%p accuracy improvement on the 3-bit quantized LLaMA2-13B and LLaMA3.2-1B models. In addition, BoA shows 20%p higher accuracy for the 3-bit quantized LLaMA3.2-3B model. The key factor leading to such an outstanding performance is that BoA considers inter-layer dependencies within the attention module by exploiting the proposed attention-aware Hessians while GPTQ assumes the independence of layers.
Comparison with transformation-based methods
To improve the quantization performance, recent studies have transformed models by applying smoothing, rotation, or permutation (Shao et al., 2023; Ma et al., 2024; Ashkboos et al., 2024; Liu et al., 2024; Lin et al., 2024a). In this experiment, we evaluate BoA under those model transformations (see Table 4 and Tables 10-11 in Section C.2). For conventional methods, we ran the official codes provided by the authors and reported the obtained results; details for implementing conventional methods are provided in Section C.2. When measuring the performance of BoA, we also transformed the model for a fair comparison. We applied QuaRot for the transformation because QuaRot does not require any training and incurs no extra costs during the actual inference (Ashkboos et al., 2024).
We observe that the performance of the proposed BoA is boosted significantly by applying the transformation. For example, the PPLs of the 2-bit quantized LLaMA2-13B / LLaMA3-8B models improve from 18.33 / 21.70 to 8.237 / 15.24, respectively (see Tables 3 and 4). For all models, we achieve at least 5%p improvement in the zero-shot accuracy (in particular 11%p improvement for the 2-bit quantized LLaMA2-13B model). We note that the performance gap between the proposed BoA and GPTQ remains significant after the model transformation; for example, the zero-shot accuracy of BoA on the 2-bit quantized LLaMA2-7B model is 12%p larger than that obtained by QuaRot-GPTQ.
Table 5: Processing time (hour) of different quantization methods considering inter-layer dependencies
Method Optimization Type LLaMA Model Size
7B 13B 30B
OmniQuant Gradient-based 1.827 3.322 7.661
AffineQuant 4.323 8.406 21.37
DuQuant + LWC 1.221 2.077 4.547
BoA One-shot 0.961 1.553 3.295
Table 6: Weight-activation quantization performance on transformed LLaMA2 and LLaMA3 models
Model Precision Method Wiki2 PPL ($\downarrow$) Zero-shot Accuracy ($\uparrow$)
Arc-c Arc-e BQ HS LAMB OBQA PIQA WG Average
LLaMA2-7B FP16 Baseline 5.473 45.90 74.66 77.92 75.94 70.86 44.00 78.89 68.90 67.13
W2A4KV4 OmniQuant 1.0e5 26.88 26.30 39.02 25.57 0.00 25.60 48.42 50.51 30.29
AffineQuant NaN NaN NaN NaN NaN NaN NaN NaN NaN NaN
SpinQuant-RTN 23.23 25.51 34.01 62.20 32.09 14.48 26.00 55.17 50.91 37.55
SpinQuant-GPTQ 24.29 22.95 36.74 59.88 32.83 13.81 26.20 56.64 51.30 37.54
DuQuant 6.4e3 28.41 26.73 38.01 25.65 0.00 26.00 48.48 51.62 30.61
DuQuant + LWC 16.35 24.91 37.33 61.99 41.46 10.54 28.20 58.71 53.28 39.55
BoA 11.80 26.79 49.20 63.09 48.05 37.76 30.80 63.55 57.85 47.14
LLaMA2-13B FP16 Baseline 4.885 49.06 77.65 80.49 79.38 73.41 45.80 80.69 72.22 69.84
W2A4KV4 OmniQuant 3.8e3 26.88 26.30 37.83 25.48 0.00 23.60 48.20 49.96 29.78
AffineQuant 1.2e3 26.62 27.19 37.83 26.37 0.00 24.60 48.80 49.17 30.07
SpinQuant-RTN 11.71 26.96 40.61 63.12 44.54 29.64 29.20 60.88 53.67 43.58
SpinQuant-GPTQ 15.54 23.55 41.29 62.17 35.76 16.50 29.80 59.52 51.54 40.02
DuQuant 6.4e3 28.41 26.73 38.01 25.65 0.00 26.00 48.48 51.62 30.61
DuQuant + LWC 16.35 24.91 37.33 61.99 41.46 10.54 28.20 58.71 53.28 39.55
BoA 8.974 31.74 55.98 64.80 56.22 49.18 34.40 68.44 59.27 52.50
LLaMA3-8B FP16 Baseline 6.137 53.67 77.61 81.19 79.15 72.23 45.00 81.01 73.24 70.39
W2A4KV4 OmniQuant 3.2e5 25.85 25.46 38.26 25.35 0.00 26.80 51.58 49.09 30.30
AffineQuant NaN NaN NaN NaN NaN NaN NaN NaN NaN NaN
SpinQuant-RTN 31.27 24.40 33.50 62.14 36.01 19.99 27.40 56.69 52.80 39.12
SpinQuant-GPTQ 29.60 22.18 34.89 58.23 35.07 13.42 27.20 55.88 51.22 37.26
DuQuant 5.4e5 27.30 24.62 50.67 26.49 0.00 28.20 50.82 50.59 32.34
DuQuant + LWC 4.1e4 25.60 26.09 37.86 25.55 0.00 25.60 51.36 49.25 30.16
BoA 18.23 28.92 48.86 62.54 44.62 29.38 28.80 62.79 54.22 45.02
LLaMA3.2-1B FP16 Baseline 13.15 38.14 63.26 69.51 60.78 54.38 34.60 74.37 59.51 56.82
W2A4KV4 OmniQuant 7.0e3 24.83 26.26 37.98 25.59 0.00 27.60 49.84 50.75 30.36
AffineQuant NaN NaN NaN NaN NaN NaN NaN NaN NaN NaN
SpinQuant-RTN 110.6 22.61 32.07 52.14 28.59 4.08 28.20 52.83 49.49 33.75
SpinQuant-GPTQ 143.8 22.35 31.23 51.96 28.91 2.91 26.20 53.32 51.85 33.59
DuQuant 5.6e4 26.62 24.71 41.74 26.16 0.01 28.20 48.80 50.51 30.84
DuQuant + LWC 8.5e3 28.07 26.01 38.26 25.84 0.00 28.80 50.11 50.91 31.00
BoA 77.05 22.61 36.66 57.22 32.13 8.07 25.80 55.55 50.99 36.13
LLaMA3.2-3B FP16 Baseline 11.04 46.16 67.80 78.62 70.44 62.15 36.00 75.52 67.40 63.01
W2A4KV4 OmniQuant 6.6e3 27.82 25.00 38.35 25.43 0.00 28.00 52.18 49.80 30.82
AffineQuant 6.0e3 25.94 26.89 38.10 25.66 0.01 28.40 51.58 49.80 30.80
SpinQuant-RTN 51.19 24.83 31.48 45.32 29.90 11.05 27.80 55.17 50.20 34.47
SpinQuant-GPTQ 65.09 23.21 33.54 38.26 31.16 7.18 25.80 55.98 51.38 33.31
DuQuant 1.2e5 26.02 25.51 49.02 26.74 0.00 29.00 51.03 49.88 32.15
DuQuant + LWC 1.7e3 24.40 26.09 37.83 25.91 0.00 27.80 50.49 52.49 30.63
BoA 37.12 27.99 37.67 59.30 37.91 15.81 27.40 57.73 52.41 39.53
  • $\dagger$
    BoA has been applied after transforming the model via SpinQuant.
    The LET option has been deactivated because this option does not support models exploiting GQA.
    ‘NaN‘ means that loss diverges in the quantization process.
    Results for other configurations are provided in Section C.3 due to the page limitation.
Finally, we emphasize that BoA not only surpasses existing transformation-based methods exploiting the naive nearest rounding (i.e., OmniQuant, AffineQuant, and DuQuant + LWC) but also is much faster than those methods relying on the gradient-based optimization (see Table 5). For example, on the LLaMA3-8B model, BoA achieves 20%p accuracy improvement for INT2 (see Table 4) and 13%p improvement for INT3 (see Table 10) over OmniQuant, AffineQuant, and DuQuant + LWC, yet reduces the quantization processing time of OmniQuant and AffineQuant more than twofold (see Table 5). Moreover, the degree of reduction increases with the model size; BoA requires 7 times shorter processing time than that needed by AffineQuant on LLaMA-30B.
4.3 Weight-Activation Quantization Results
We now evaluate the weight-activation quantization performance (see Table 6 and Tables 12-14 in Section C.3). As in previous studies (Shao et al., 2023; Ma et al., 2024; Lin et al., 2024a; Ashkboos et al., 2024; Liu et al., 2024), we apply per-token nearest-rounding quantization for input activations and KV caches. We use the notation ‘WxAyKVz’ to denote the x-bit weight quantization, y-bit input activation quantization, and z-bit KV cache quantization. When measuring the performance of RTN, GPTQ, and the proposed BoA, we use SpinQuant for the model transformation as SpinQuant outperforms QuaRot by optimizing the transformation for the activation quantization (Liu et al., 2024).
From Table 6 and Tables 12-14 in Section C.3, we observe that the outstanding weight-quantization performance of BoA leads to state-of-the-art performance for the weight-activation quantization. In particular, the performance gain is noticeable for low-bit (e.g., W2A4KV4 and W2A4KV16). For example, BoA achieves 10%p accuracy improvement on LLaMA2-7B and 12.5%p improvement on LLaMA2-13B over SpinQuant-GPTQ (see Table 6). Furthermore, compared to OmniQuant, AffineQuant, and DuQuant, BoA achieves 8%p, 13%p, 13%p, 5%p, and 7%p improvement on LLaMA2-7B, LLaMA2-13B, LLaMA3-8B, LLaMA3.2-1B, and LLaMA3.2-3B, respectively.
5 Conclusion
In this paper, we proposed a novel backpropagation-free PTQ algorithm called BoA. To consider the inter-layer dependencies within the attention module, we approximated the Hessian matrices by exploiting the attention reconstruction error rather than the layer-wise reconstruction error. Furthermore, to mitigate the computational overhead incurred by the proposed attention-aware Hessians, we incorporated several techniques including Hessian relaxation, efficient computation of inverse Hessians, and head-wise simultaneous quantization. Finally, from extensive experiments, we demonstrated the efficacy of the proposed BoA.
Impact Statement
BoA is a highly efficient and accurate quantization algorithm that minimizes accuracy loss while significantly reducing the time required for model quantization and deployment. By synergizing with other methods, BoA facilitates the efficient operation of LLMs on resource-constrained hardware using only integer arithmetic. This breakthrough is particularly impactful for on-device AI, allowing real-time model inference on mobile and edge devices without the need for server access, paving the way for scalable, decentralized AI solutions.
Appendix A Proof of Footnote 2
In our proof, we use the following useful properties of the Kronecker product:
$\operatornamewithlimits{vec}\left(\mathbf{M}_{1}\mathbf{M}_{2}\mathbf{M}_{3}\right)$ $=\left(\mathbf{M}_{3}^{T}\otimes\mathbf{M}_{1}\right)\operatornamewithlimits{vec}(\mathbf{M}_{2}),$ (13a)
$\left(\mathbf{M}_{1}\otimes\mathbf{M}_{2}\right)^{T}$ $=\mathbf{M}_{1}^{T}\otimes\mathbf{M}_{2}^{T},$ (13b)
$\left(\mathbf{M}_{1}\otimes\mathbf{M}_{2}\right)\left(\mathbf{M}_{3}\otimes\mathbf{M}_{4}\right)$ $=\mathbf{M}_{1}\mathbf{M}_{3}\otimes\mathbf{M}_{2}\mathbf{M}_{4},$ (13c)
where $\operatornamewithlimits{vec}(\cdot)$ denotes the vectorization operation.
Using 13a, we have
$\left\|\mathbf{M}_{1}\Delta\mathbf{W}\mathbf{M}_{2}\right\|_{F}^{2}$ $=\left\|\left(\mathbf{M}_{2}^{T}\otimes\mathbf{M}_{1}\right)\Delta\mathbf{w}\right\|_{2}^{2}=\Delta\mathbf{w}^{T}\left(\mathbf{M}_{2}^{T}\otimes\mathbf{M}_{1}\right)^{T}\left(\mathbf{M}_{2}^{T}\otimes\mathbf{M}_{1}\right)\Delta\mathbf{w},$
where $\Delta\mathbf{w}=\operatornamewithlimits{vec}(\Delta\mathbf{W})$. In addition, by 13b and 13c, we have
$\Delta\mathbf{w}^{T}\left(\mathbf{M}_{2}^{T}\otimes\mathbf{M}_{1}\right)^{T}\left(\mathbf{M}_{2}^{T}\otimes\mathbf{M}_{1}\right)\Delta\mathbf{w}$ $=\Delta\mathbf{w}^{T}\left(\mathbf{M}_{2}\otimes\mathbf{M}_{1}^{T}\right)\left(\mathbf{M}_{2}^{T}\otimes\mathbf{M}_{1}\right)\Delta\mathbf{w}$
$=\Delta\mathbf{w}^{T}\left(\mathbf{M}_{2}\mathbf{M}_{2}^{T}\otimes\mathbf{M}_{1}^{T}\mathbf{M}_{1}\right)\Delta\mathbf{w}.$
Finally, by exploiting the fact that $\frac{\partial^{2}\mathbf{x}^{T}\mathbf{A}\mathbf{x}}{\partial\mathbf{x}^{2}}=\mathbf{A}+\mathbf{A}^{T}$, we obtain
$\frac{\partial^{2}\left\|\mathbf{M}_{1}\Delta\mathbf{W}\mathbf{M}_{2}\right\|_{F}^{2}}{\partial\Delta\mathbf{w}^{2}}$ $=\mathbf{M}_{2}\mathbf{M}_{2}^{T}\otimes\mathbf{M}_{1}^{T}\mathbf{M}_{1}+\left(\mathbf{M}_{2}\mathbf{M}_{2}^{T}\otimes\mathbf{M}_{1}^{T}\mathbf{M}_{1}\right)^{T}$
=$\overset{(a)}{=}\mathbf{M}_{2}\mathbf{M}_{2}^{T}\otimes\mathbf{M}_{1}^{T}\mathbf{M}_{1}+\left(\mathbf{M}_{2}\mathbf{M}_{2}^{T}\right)^{T}\otimes\left(\mathbf{M}_{1}^{T}\mathbf{M}_{1}\right)^{T}$
$=2\mathbf{M}_{2}\mathbf{M}_{2}^{T}\otimes\mathbf{M}_{1}^{T}\mathbf{M}_{1},$
where (a) follows from 13b. This completes the proof.
Appendix B Refined Weight-update Formula
We recall that the Hessian-based weight-update formula is given by (Frantar & Alistarh, 2022; Frantar et al., 2023)
$\bm{\delta}{\bm{w}}$ $=-\frac{w_{q}-{\mathcal{Q}}(w_{q})}{[\mathbf{U}]_{q,q}}[\mathbf{U}]_{q,:}\text{ where }\mathbf{U}=\operatornamewithlimits{Chol}(\mathbf{H}^{-1})^{T}.$
For the proposed attention-aware Hessians in Table 1, we have
$\mathbf{U}_{h}=\mathbf{U}_{\text{col},h}\otimes\mathbf{U}_{\text{row},h},$
where $\mathbf{U}_{\text{col},h}=\operatornamewithlimits{Chol}(\mathbf{H}_{\text{col},h}^{-1})^{T}$ and $\mathbf{U}_{\text{row},h}=\operatornamewithlimits{Chol}(\mathbf{H}_{\text{row},h}^{-1})^{T}$ (see Section 3.3). Therefore, the weight-update formula can be recast as
$\bm{\delta}{\bm{w}}_{h}$ $=-\frac{w_{q}-{\mathcal{Q}}(w_{q})}{[\mathbf{U}_{\text{col},h}\otimes\mathbf{U}_{\text{row},h}]_{q,q}}[\mathbf{U}_{\text{col},h}\otimes\mathbf{U}_{\text{row},h}]_{q,:}.$
Figure 2: Illustration of the Hessian information when $d_{\text{row}}=2$ and $d_{\text{col}}=3$
For simplicity, suppose we quantize the first (00-th) row. When the weight $[\mathbf{W}_{h}]_{0,j}(=[\mathbf{W}^{(0)}]_{h,j})$ in the $j$-th column is quantized, the weight-update of the $i$-th row is simplified as (see Figure 2 for the ease of understanding)
$[\bm{\delta}\mathbf{W}_{h}]_{i,:}$ $=-\frac{[\mathbf{W}_{h}]_{0,j}-{\mathcal{Q}}([\mathbf{W}_{h}]_{0,j})}{[\mathbf{U}_{\text{row},h}]_{0,0}[\mathbf{U}_{\text{col},h}]_{j,j}}[\mathbf{U}_{\text{row},h}]_{0,i}[\mathbf{U}_{\text{col},h}]_{j,:}$
$=-\frac{[\mathbf{W}_{h}]_{0,j}-{\mathcal{Q}}([\mathbf{W}_{h}]_{0,j})}{[\mathbf{U}_{\text{col},h}]_{j,j}}\cdot\frac{[\mathbf{U}_{\text{row},h}]_{0,i}[\mathbf{U}_{\text{col},h}]_{j,:}}{[\mathbf{U}_{\text{row},h}]_{0,0}}$
Thus, after the quantization of all weights in the first row, the total amount of the weight-update for the $i$-th row can be expressed as
$[\bm{\delta}\mathbf{W}_{h,\text{total}}]_{i,:}$ $=-\sum_{j=0}^{d_{\text{col}}-1}\frac{[\mathbf{W}_{h}]_{0,j}-{\mathcal{Q}}([\mathbf{W}_{h}]_{0,j})}{[\mathbf{U}_{\text{col},h}]_{j,j}}\cdot\frac{[\mathbf{U}_{\text{row},h}]_{0,i}[\mathbf{U}_{\text{col},h}]_{j,:}}{[\mathbf{U}_{\text{row},h}]_{0,0}}$
$=-\frac{[\mathbf{U}_{\text{row},h}]_{0,i}}{[\mathbf{U}_{\text{row},h}]_{0,0}}\sum_{j=0}^{d_{\text{col}}-1}\frac{[\mathbf{W}_{h}]_{0,j}-{\mathcal{Q}}([\mathbf{W}_{h}]_{0,j})}{[\mathbf{U}_{\text{col},h}]_{j,j}}\cdot[\mathbf{U}_{\text{col},h}]_{j,:}.$
Furthermore, by noting that (see line 5 in Algorithm 2)
$[\mathbf{E}_{\text{GPTQ}}]_{h,j}$ $=\frac{[\mathbf{W}_{h}]_{0,j}-{\mathcal{Q}}([\mathbf{W}_{h}]_{0,j})}{[\mathbf{U}_{\text{col},h}]_{j,j}},$
we obtain
$[\bm{\delta}\mathbf{W}_{h,\text{total}}]_{i,:}$ $=-\frac{[\mathbf{U}_{\text{row},h}]_{0,i}}{[\mathbf{U}_{\text{row},h}]_{0,0}}\sum_{j=0}^{d_{\text{col}}-1}[\mathbf{E}_{\text{GPTQ}}]_{h,j}\cdot[\mathbf{U}_{\text{col},h}]_{j,:}=-\frac{[\mathbf{U}_{\text{row},h}]_{0,i}}{[\mathbf{U}_{\text{row},h}]_{0,0}}[\mathbf{E}_{\text{GPTQ}}]_{h,:}\mathbf{U}_{\text{col},h}.$
As a result, the weight-update matrix to compensate for the quantization error of the first row is given by
$[\bm{\delta}\mathbf{W}_{h,\text{total}}]_{0:,:}$ $=-\frac{[\mathbf{U}_{\text{row},h}^{T}]_{0:,0}[\mathbf{E}_{\text{GPTQ}}]_{h,:}\mathbf{U}_{\text{col},h}}{[\mathbf{U}_{\text{row},h}]_{0,0}}.$ (14)
By taking similar steps as above, we can easily generalize 14 for the $j$-th row as follows:
$[\bm{\delta}\mathbf{W}_{h,\text{total}}]_{j:,:}$ $=-\frac{[\mathbf{U}_{\text{row},h}^{T}]_{j:,j}[\mathbf{E}_{\text{GPTQ}}]_{h,:}\mathbf{U}_{\text{col},h}}{[\mathbf{U}_{\text{row},h}]_{j,j}}.$ (15)
Appendix C Additional Experimental Results
C.1 Performance of Weight-Rounding Optimization
We evaluate the weight-only quantization performance of the proposed BoA. To solely compare the performance of the weight-rounding optimization, we do not apply any transform (e.g., smoothing, rotation, and permutation) to the models. In this appendix, we provide the results for the INT4 weight quantization of LLaMA2 and LLaMA3 models (see Table 7) and the results on LLaMA1 models (see Table 8) and OPT models (see Table 9).
Table 7: INT4 weight-only quantization performance on LLaMA2 and LLaMA3 models without transformation
Model Precision Method Wiki2 PPL ($\downarrow$) Zero-shot Accuracy ($\uparrow$)
Arc-c Arc-e BQ HS LAMB OBQA PIQA WG Average
LLaMA2-7B FP Baseline 5.473 45.90 74.66 77.92 75.94 70.86 44.00 78.89 68.90 67.13
INT4 RTN 6.998 41.64 67.05 72.72 71.36 59.78 42.60 77.04 65.75 62.24
GPTQ 5.809 42.06 69.82 67.06 70.44 62.80 43.00 77.15 68.11 62.56
BoA 5.622 43.94 72.26 78.04 74.53 68.96 44.20 78.40 69.14 66.18
LLaMA2-13B FP Baseline 4.885 49.06 77.65 80.49 79.38 73.41 45.80 80.69 72.22 69.84
INT4 RTN 6.319 43.00 66.88 71.01 66.14 62.25 38.20 77.04 63.85 61.05
GPTQ 5.632 47.18 74.49 74.37 75.00 70.89 44.20 78.40 69.69 66.78
BoA 5.096 47.95 75.80 78.93 77.96 70.61 41.60 79.87 71.35 68.01
LLaMA3-8B FP Baseline 6.137 53.67 77.61 81.19 79.15 72.23 45.00 81.01 73.24 70.39
INT4 RTN 7.981 49.06 73.65 73.30 76.62 59.55 45.00 78.29 72.93 66.05
GPTQ 7.960 49.06 73.44 73.70 69.30 60.19 45.00 79.54 73.56 65.47
BoA 6.561 50.17 77.74 80.31 77.42 70.20 44.40 80.14 73.88 69.28
LLaMA3.2-1B FP Baseline 13.15 38.14 63.26 69.51 60.78 54.38 34.60 74.37 59.51 56.82
INT4 RTN 18.81 34.30 57.41 65.20 55.93 32.72 33.00 69.15 56.35 50.51
GPTQ 16.20 35.41 58.63 64.46 55.69 38.08 32.20 69.04 56.27 51.22
BoA 14.29 36.18 60.82 67.22 59.03 51.65 34.00 71.87 58.09 54.86
LLaMA3.2-3B FP Baseline 11.04 46.16 67.80 78.62 70.44 62.15 36.00 75.52 67.40 63.01
INT4 RTN 13.86 43.26 61.24 73.67 68.27 54.18 38.40 71.76 64.01 59.35
GPTQ 12.41 44.28 63.51 77.00 68.44 57.39 36.60 74.16 67.01 61.05
BoA 11.74 44.11 66.75 77.86 69.42 61.74 36.20 75.08 66.85 62.25
Table 8: Weight-only quantization performance on LLaMA models without transformation
Model Precision Method Wiki2 PPL ($\downarrow$) Zero-shot Accuracy ($\uparrow$)
Arc-c Arc-e BQ HS LAMB OBQA PIQA WG Average
LLaMA-7B FP Baseline 5.677 44.71 72.94 75.05 76.21 70.66 44.40 79.16 70.01 66.64
INT2 RTN 5.9e3 26.37 26.73 39.82 26.37 0.01 26.00 51.80 48.62 30.72
GPTQ 19.63 25.34 41.58 56.27 37.27 26.24 30.20 61.70 55.56 41.77
BoA 10.28 28.75 52.69 64.19 49.34 41.48 32.60 66.54 59.19 49.35
INT3 RTN 61.31 26.45 50.46 59.76 40.07 14.64 33.00 67.08 55.88 43.42
GPTQ 18.44 33.53 61.36 64.40 56.22 44.91 35.40 72.09 64.33 54.03
BoA 6.267 39.76 67.68 75.35 71.01 68.92 41.80 77.80 67.48 63.73
INT4 RTN 7.912 41.47 70.29 72.81 73.16 63.81 41.60 77.86 68.27 63.66
GPTQ 7.791 40.36 66.50 70.58 68.48 61.84 39.00 77.42 68.43 61.58
BoA 5.899 43.77 72.18 75.54 74.75 69.60 44.40 78.62 69.77 66.08
LLaMA-13B FP Baseline 5.090 47.70 74.75 77.95 79.07 73.66 44.80 80.14 72.77 68.86
INT2 RTN 2.2e3 25.60 26.01 39.48 26.00 0.00 25.20 49.46 48.54 30.04
GPTQ 34.63 26.37 43.43 56.67 46.85 33.11 31.40 65.07 57.62 45.07
BoA 8.309 32.76 59.51 69.57 59.52 51.79 35.20 71.22 65.19 55.60
INT3 RTN 134.5 25.00 49.24 56.39 36.91 11.41 29.20 64.20 54.62 40.87
GPTQ 7.096 42.24 69.74 70.73 73.40 62.99 39.20 77.09 68.11 62.94
BoA 5.461 45.82 71.93 75.96 75.71 70.62 43.40 78.35 70.24 66.50
INT4 RTN 7.742 45.22 69.99 72.42 75.67 65.65 41.60 78.78 70.56 64.99
GPTQ 6.147 45.82 72.26 74.10 76.24 70.15 43.80 80.20 71.90 66.81
BoA 5.189 47.44 74.62 77.31 78.03 73.19 45.40 80.20 73.09 68.66
LLaMA-30B FP Baseline 4.101 52.90 78.96 82.69 82.63 75.47 48.20 82.26 75.77 72.36
INT2 RTN 7.1e3 25.60 26.52 40.67 27.20 0.11 24.60 50.92 49.64 30.66
GPTQ 9.770 35.41 60.14 64.77 61.54 50.12 38.20 70.35 66.54 55.88
BoA 6.669 37.71 64.73 73.00 64.77 61.28 39.00 73.34 67.40 60.15
INT3 RTN 81.76 30.12 58.21 54.10 42.94 6.71 35.00 69.31 58.25 44.33
GPTQ 7.064 48.12 71.46 74.89 76.73 66.39 43.20 78.35 70.80 66.24
BoA 4.575 49.57 74.75 81.87 79.21 73.77 46.20 79.38 74.19 69.87
INT4 RTN 5.802 49.40 74.16 78.07 79.48 66.67 44.00 79.87 73.16 68.10
GPTQ 5.188 51.45 75.67 80.18 79.69 70.38 45.40 79.43 73.64 69.48
BoA 4.218 53.16 78.41 82.72 81.95 75.96 49.40 81.88 75.06 72.32
Table 9: Weight-only quantization performance on OPT models without transformation
Model Precision Method Wiki2 PPL ($\downarrow$) Zero-shot Accuracy ($\uparrow$)
Arc-c Arc-e BQ HS LAMB OBQA PIQA WG Average
OPT-125M FP Baseline 27.65 22.78 39.98 55.44 31.35 33.44 28.00 62.02 50.28 40.41
INT2 RTN 1.0e4 23.89 27.31 37.83 25.96 0.00 27.00 50.49 51.14 30.45
GPTQ 411.3 22.61 31.52 38.47 27.16 0.18 27.20 52.18 51.70 31.38
BoA 85.63 21.42 32.15 48.20 28.50 6.86 26.60 56.47 49.17 33.67
INT3 RTN 233.9 22.01 33.33 52.94 28.52 2.81 28.40 59.19 51.14 34.79
GPTQ 50.75 21.67 33.84 58.38 29.41 16.96 25.40 58.98 51.14 36.97
BoA 31.95 22.95 36.83 59.66 30.67 27.10 28.80 60.34 50.83 39.65
OPT-350M FP Baseline 22.00 23.89 40.36 57.68 36.67 40.47 28.20 64.74 52.33 43.04
INT2 RTN 4.6e3 24.74 27.65 37.83 26.87 0.00 23.40 52.83 52.09 30.68
GPTQ 61.93 22.53 33.46 57.28 29.66 10.70 23.80 56.75 51.22 35.68
BoA 46.53 22.01 35.90 62.32 30.43 11.85 26.60 58.98 51.62 37.46
INT3 RTN 35.91 23.38 36.99 52.14 34.00 30.37 27.00 62.02 51.62 39.69
GPTQ 25.25 22.95 38.47 60.58 35.25 41.38 27.60 62.46 51.22 42.49
BoA 23.96 22.87 38.76 61.22 35.30 38.08 28.40 62.57 53.28 42.56
OPT-1.3B FP Baseline 14.62 29.52 50.97 57.83 53.70 55.19 33.40 72.47 59.51 51.57
INT2 RTN 1.6e4 24.23 24.75 51.19 26.51 0.00 29.00 52.29 50.04 32.25
GPTQ 177.4 22.78 31.78 57.06 28.13 6.37 24.80 53.75 48.70 34.17
BoA 31.65 22.70 37.67 62.17 34.65 18.02 26.80 61.48 51.78 39.41
INT3 RTN 754.8 25.34 34.72 52.17 36.31 10.25 27.20 60.07 51.78 37.23
GPTQ 19.35 27.13 43.43 54.25 47.23 36.30 29.40 68.01 54.30 45.01
BoA 15.77 27.39 48.74 58.99 50.09 51.50 30.60 69.37 57.30 49.25
OPT-2.7B FP Baseline 12.47 31.31 54.38 60.40 60.62 59.78 35.20 74.81 61.01 54.69
INT2 RTN 2.7e5 26.11 26.73 62.17 26.77 0.00 29.20 50.87 51.07 34.12
GPTQ 135.5 24.66 33.71 60.28 33.62 17.62 27.00 57.73 50.59 38.15
BoA 22.30 25.77 41.62 62.26 42.00 33.76 28.80 62.08 52.96 43.66
INT3 RTN 170.5 25.26 38.34 56.51 36.64 14.61 29.80 66.27 53.04 40.06
GPTQ 15.64 29.27 49.07 45.90 51.79 42.92 33.00 69.80 57.54 47.41
BoA 13.63 29.10 52.90 65.29 55.86 56.94 34.00 72.03 60.06 53.27
OPT-6.7B FP Baseline 10.86 34.73 60.02 66.06 67.16 65.50 37.40 76.50 65.43 59.10
INT2 RTN 2.9e4 25.00 25.38 37.83 26.66 0.00 29.20 50.00 51.46 30.69
GPTQ 95.95 23.21 33.75 61.77 32.09 16.25 25.80 57.02 52.01 37.74
BoA 20.57 28.24 47.31 61.80 47.00 34.09 33.20 66.87 57.54 47.01
INT3 RTN 53.76 27.73 40.87 50.95 48.67 28.32 31.20 68.17 53.12 43.63
GPTQ 12.05 32.51 56.40 54.53 62.49 55.99 36.60 74.21 60.85 54.20
BoA 11.25 32.76 57.32 66.85 64.53 63.64 37.60 75.57 64.09 57.80
OPT-13B FP Baseline 10.13 35.75 61.87 65.84 69.87 65.98 39.00 76.88 65.11 60.04
INT2 RTN 4.7e4 26.79 26.64 40.40 26.25 0.00 27.40 49.78 50.04 30.91
GPTQ 49.40 26.28 37.04 62.14 41.60 29.17 27.60 61.04 52.17 42.13
BoA 15.34 29.27 48.95 62.78 53.63 44.71 32.60 68.50 58.64 49.89
INT3 RTN 40.46 27.47 38.72 62.54 51.21 29.44 26.80 63.38 49.88 43.68
GPTQ 11.11 34.56 59.51 62.45 64.81 58.57 36.00 74.92 63.54 56.80
BoA 10.39 34.30 59.93 69.94 66.95 66.61 37.60 75.95 66.93 59.78
OPT-30B FP Baseline 9.557 38.05 65.36 70.46 72.27 67.85 40.20 78.07 68.43 62.59
INT2 RTN 4.2e4 26.79 25.55 37.83 25.88 0.00 28.60 50.60 51.22 30.81
GPTQ 21.19 27.99 41.96 62.45 46.90 39.48 30.40 65.72 54.22 46.14
BoA 11.15 30.55 55.81 62.81 60.19 60.39 34.80 72.47 62.67 54.96
INT3 RTN 86.63 24.32 35.06 57.98 38.07 12.17 26.00 59.90 52.09 38.20
GPTQ 10.14 34.90 61.70 66.97 69.34 63.70 40.40 77.26 67.01 60.16
BoA 9.744 36.26 62.63 70.28 70.52 68.86 40.20 77.75 67.17 61.71
C.2 Comparison with Transformation-based Methods
We compare the weight-only quantization performances of the proposed BoA and existing transformation-based methods. For conventional methods, we ran the official codes provided by the authors and reported the obtained results. For OmniQuant (Shao et al., 2023) and AffineQuant (Ma et al., 2024), we activated both learnable equivalent transformation (LET) and learnable weight clipping (LWC) options. For AffineQuant, we did not activate the use-ln-matrix option because this adds extra affine transformation between the normalization layer and linear layers, which incurs additional processing time during the inference. QuaRot-RTN and QuaRot-GPTQ mean that RTN and GPTQ have been applied for the weight quantization after transforming models via QuaRot (Ashkboos et al., 2024). For DuQuant, we report the results obtained with and without activating the LWC option as in the original paper (Lin et al., 2024a). When measuring the performance of the proposed BoA, we also transformed the model for fair comparison. We applied QuaRot for the transformation because QuaRot does not require training and incurs no extra costs during the actual inference (Ashkboos et al., 2024).
In this appendix, we provide the results for the INT3 weight quantization of LLaMA2 and LLaMA3 models (see Table 10) and the results on LLaMA1 models (see Table 11).
Table 10: INT3 weight-only quantization performance on transformed LLaMA2 and LLaMA3 models
Model Precision Method Wiki2 PPL ($\downarrow$) Zero-shot Accuracy ($\uparrow$)
Arc-c Arc-e BQ HS LAMB OBQA PIQA WG Average
LLaMA2-7B FP16 Baseline 5.473 45.90 74.66 77.92 75.94 70.86 44.00 78.89 68.90 67.13
INT3 OmniQuant 6.640 39.59 65.66 69.82 70.09 57.58 38.40 76.01 64.88 60.25
AffineQuant 6.468 40.36 68.01 70.86 70.32 61.87 38.60 76.12 65.27 61.43
QuaRot-RTN 129.2 23.46 30.60 37.83 29.30 4.75 24.40 53.10 50.20 31.71
QuaRot-GPTQ 6.122 40.96 70.16 73.30 71.44 67.63 41.40 77.42 67.48 63.72
DuQuant 6.831 41.13 67.59 69.36 70.52 62.25 41.40 76.01 65.35 61.70
DuQuant + LWC 6.226 40.70 69.15 74.77 70.97 64.06 42.40 77.04 66.85 63.24
BoA 5.874 42.06 71.17 73.73 72.29 69.85 41.60 77.37 67.48 64.44
LLaMA2-13B FP16 Baseline 4.885 49.06 77.65 80.49 79.38 73.41 45.80 80.69 72.22 69.84
INT3 OmniQuant 5.593 44.80 71.72 75.29 75.63 64.94 43.20 78.67 69.30 65.44
AffineQuant 5.526 45.73 73.74 77.34 74.79 65.56 43.00 77.15 67.01 65.54
QuaRot-RTN 48.06 22.01 34.72 62.14 33.44 7.80 26.60 57.67 50.51 36.86
QuaRot-GPTQ 5.382 47.01 75.42 78.84 75.70 72.14 44.60 78.56 70.01 67.79
DuQuant 5.749 44.88 72.10 79.91 75.37 68.23 43.00 77.80 70.72 66.50
DuQuant + LWC 5.414 46.67 74.62 77.92 75.93 68.63 43.80 79.98 68.75 67.04
BoA 5.202 48.29 76.39 79.45 76.23 73.19 45.20 78.67 70.96 68.55
LLaMA3-8B FP16 Baseline 6.137 53.67 77.61 81.19 79.15 72.23 45.00 81.01 73.24 70.39
INT3 OmniQuant 11.75 37.37 60.94 65.57 66.81 33.02 34.80 71.49 59.59 53.70
AffineQuant 11.63 37.97 63.64 64.92 67.42 33.95 36.20 71.98 60.30 54.55
QuaRot-RTN 38.64 23.72 38.22 51.28 47.53 23.63 33.60 62.35 62.04 42.80
QuaRot-GPTQ 7.490 47.87 74.49 78.44 74.67 67.43 40.00 78.07 72.38 66.67
DuQuant 11.35 35.84 53.37 64.56 65.56 49.52 36.20 72.25 65.90 55.40
DuQuant + LWC 10.78 31.66 46.59 64.65 65.25 41.29 38.00 70.13 61.80 52.42
BoA 7.145 49.06 77.40 80.49 74.54 69.20 43.00 78.51 72.53 68.09
LLaMA3.2-1B FP16 Baseline 13.15 38.14 63.26 69.51 60.78 54.38 34.60 74.37 59.51 56.82
INT3 OmniQuant 26.61 31.31 51.39 63.67 49.18 18.83 29.60 67.57 55.49 45.88
AffineQuant 26.43 32.00 52.48 64.31 48.71 20.67 26.20 67.57 52.88 45.60
QuaRot-RTN 98.24 24.06 36.03 61.96 35.31 7.68 28.60 59.14 51.62 38.05
QuaRot-GPTQ 16.56 32.42 57.95 64.10 52.48 43.79 32.20 68.99 57.14 51.13
DuQuant 2.1e4 25.26 25.67 41.35 26.68 0.00 27.80 49.51 48.78 30.63
DuQuant + LWC 2.7e4 25.68 26.73 40.73 25.34 0.00 29.00 49.40 52.09 31.12
BoA 15.73 32.42 57.95 66.24 53.86 48.74 33.20 70.18 57.06 52.46
LLaMA3.2-3B FP16 Baseline 11.04 46.16 67.80 78.62 70.44 62.15 36.00 75.52 67.40 63.01
INT3 OmniQuant 16.97 36.01 57.58 71.47 59.57 39.17 34.00 69.31 60.22 53.42
AffineQuant 16.79 36.43 56.57 72.97 59.19 40.07 35.20 70.13 60.06 53.83
QuaRot-RTN 89.54 22.01 32.79 59.08 36.18 4.82 25.00 54.35 53.67 35.99
QuaRot-GPTQ 13.58 38.31 58.96 75.72 64.59 54.59 35.80 70.18 64.96 57.89
DuQuant 18.92 30.55 48.32 71.71 60.34 40.54 32.40 66.54 60.38 51.35
DuQuant + LWC 15.18 37.20 58.33 72.02 61.22 44.58 33.40 70.02 59.98 54.59
BoA 12.97 42.49 65.74 78.90 65.59 58.40 34.80 73.07 63.46 60.31
  • $\dagger$
    BoA has been applied after transforming the model via QuaRot.
    The LET option has been deactivated because this option does not support models exploiting GQA.
Table 11: Weight-only quantization performance on transformed LLaMA models
Model Precision Method Wiki2 PPL ($\downarrow$) Zero-shot Accuracy ($\uparrow$)
Arc-c Arc-e BQ HS LAMB OBQA PIQA WG Average
LLaMA-7B FP Baseline 5.677 44.71 72.94 75.05 76.21 70.66 44.40 79.16 70.01 66.64
INT2 OmniQuant 15.74 26.28 45.71 61.93 42.66 17.43 29.00 62.40 52.17 42.20
AffineQuant 11.93 27.22 49.33 62.51 49.26 23.57 32.20 62.89 54.06 45.13
QuaRot-RTN 1.0e4 27.90 26.39 37.83 26.03 0.00 21.60 49.24 49.49 29.81
QuaRot-GPTQ 11.53 26.45 45.75 62.48 48.23 37.37 32.80 65.67 57.46 47.03
DuQuant 9.4e3 28.33 26.89 46.42 26.24 0.00 21.20 50.71 51.62 31.43
DuQuant + LWC 12.39 27.73 49.37 63.03 49.88 23.65 31.80 64.96 56.75 45.90
BoA 9.812 30.63 54.88 63.55 52.20 46.50 33.00 68.72 61.01 51.31
INT3 OmniQuant 6.463 38.74 66.96 72.35 69.84 62.79 40.00 77.20 66.69 61.82
AffineQuant 6.392 39.68 66.46 72.87 70.52 64.99 39.40 77.42 65.04 62.05
QuaRot-RTN 24.03 29.86 53.11 61.13 50.30 24.53 31.40 69.37 57.30 47.13
QuaRot-GPTQ 6.166 43.09 70.37 72.02 72.18 65.34 43.20 78.29 68.19 64.09
DuQuant 6.976 37.37 64.23 71.13 70.81 60.45 39.20 76.39 65.98 60.70
DuQuant + LWC 6.328 38.74 66.20 73.73 71.24 63.78 42.40 77.31 68.51 62.74
BoA 6.091 41.04 68.18 71.77 72.26 69.32 40.40 78.18 68.67 63.73
LLaMA-13B FP Baseline 5.090 47.70 74.75 77.95 79.07 73.66 44.80 80.14 72.77 68.86
INT2 OmniQuant 13.29 29.69 53.79 62.29 55.31 19.52 31.60 70.29 57.14 47.45
AffineQuant 10.65 28.75 53.54 65.57 53.03 30.73 34.00 68.12 59.35 49.14
QuaRot-RTN 4.1e3 28.16 26.47 37.83 25.72 0.00 22.60 51.25 48.46 30.06
QuaRot-GPTQ 9.331 31.06 58.59 63.79 54.96 45.54 35.20 69.97 63.30 52.80
DuQuant 6.2e3 25.94 25.93 39.66 26.70 0.00 23.60 50.87 51.62 30.54
DuQuant + LWC 8.770 31.74 58.38 66.27 60.98 44.46 37.00 72.20 62.19 54.15
BoA 8.341 33.19 63.64 71.44 60.75 56.30 37.00 71.93 65.90 57.52
INT3 OmniQuant 5.671 45.05 71.13 75.44 75.35 66.82 44.40 78.56 69.46 65.78
AffineQuant 5.615 43.00 70.79 74.83 74.99 68.95 43.40 79.76 69.38 65.64
QuaRot-RTN 7.082 37.37 60.14 73.18 67.93 56.42 35.60 76.22 66.14 59.13
QuaRot-GPTQ 5.471 44.37 71.63 76.94 75.59 72.45 43.00 78.02 71.51 66.69
DuQuant 5.923 43.94 70.12 74.40 75.76 68.33 42.00 78.56 72.06 65.65
DuQuant + LWC 5.554 45.31 71.97 72.78 75.85 69.81 45.00 79.49 70.64 66.36
BoA 5.411 45.56 72.81 75.60 75.95 72.71 45.20 79.27 71.03 67.27
LLaMA-30B FP Baseline 4.101 52.90 78.96 82.69 82.63 75.47 48.20 82.26 75.77 72.36
INT2 OmniQuant 8.598 33.45 57.37 62.84 59.21 39.36 37.20 70.02 60.14 52.45
AffineQuant 7.267 38.14 65.07 73.33 67.61 53.75 38.60 74.21 64.96 59.46
QuaRot-RTN 3.8e3 25.77 26.52 37.83 25.80 0.01 23.80 51.69 47.83 29.91
QuaRot-GPTQ 7.283 39.08 68.01 65.44 65.54 58.15 38.40 73.39 67.56 59.45
DuQuant 5.7e3 27.13 26.98 38.75 26.76 0.00 26.20 49.89 49.17 30.61
DuQuant + LWC 7.706 38.48 64.60 70.00 67.56 48.08 36.40 72.96 65.11 57.90
BoA 6.525 41.47 68.01 64.65 67.63 66.31 41.00 74.21 70.32 61.70
INT3 OmniQuant 4.766 50.09 76.18 79.91 79.58 70.79 45.20 79.92 74.43 69.51
AffineQuant 4.729 50.60 76.81 79.85 79.60 72.83 44.60 80.52 74.74 69.94
QuaRot-RTN 6.355 37.80 65.32 67.98 64.79 47.46 38.80 72.80 67.17 57.77
QuaRot-GPTQ 4.759 49.23 77.19 81.22 79.97 75.98 45.20 80.41 75.61 70.60
DuQuant 5.066 49.49 74.71 79.97 79.87 72.78 46.40 79.54 74.51 69.66
DuQuant + LWC 4.634 49.32 75.17 81.35 80.26 72.29 46.20 81.01 73.72 69.92
BoA 4.602 52.73 77.23 80.83 80.04 75.91 45.60 80.79 76.16 71.16
  • $\dagger$
    BoA has been applied after transforming the model via QuaRot.
C.3 Weight-Activation Quantization Results
We evaluate the weight-activation quantization performances of the proposed BoA. As in prior works (Shao et al., 2023; Ma et al., 2024; Lin et al., 2024a; Ashkboos et al., 2024; Liu et al., 2024), we apply per-token nearest-rounding quantization for input activations and KV caches. We use the notation ‘WxAyKVz’ to denote the x-bit weight quantization, y-bit input activation quantization, and z-bit KV cache quantization. When measuring the performance of RTN, GPTQ, and the proposed BoA, we use SpinQuant for the model transformation as SpinQuant outperforms QuaRot by optimizing the transformation for the activation quantization (Liu et al., 2024). When optimizing the rotation matrix in SpinQuant, we quantize both weights and activations for SpinQuant-RTN and quantize only activations for SpinQuant-GPTQ and the proposed BoA, as in the original paper (Liu et al., 2024). In this appendix, we provide the results on LLaMA3 models for various quantization configurations.
Table 12: Weight-activation quantization performance on the transformed LLaMA3-8B
Precision Method Wiki2 PPL ($\downarrow$) Zero-shot Accuracy ($\uparrow$)
Arc-c Arc-e BQ HS LAMB OBQA PIQA WG Average
FP16 Baseline 6.137 53.67 77.61 81.19 79.15 72.23 45.00 81.01 73.24 70.39
W2A4KV16 OmniQuant 2.9e5 25.94 26.22 38.50 24.86 0.00 28.00 50.44 49.33 30.41
AffineQuant 3.2e5 26.62 26.30 38.13 25.62 0.00 29.60 49.51 52.57 31.04
SpinQuant-RTN 28.38 25.77 35.52 62.11 37.55 18.89 24.20 56.42 51.85 39.04
SpinQuant-GPTQ 26.35 23.38 38.30 60.43 37.34 16.44 26.20 58.71 53.51 39.29
DuQuant 4.5e5 25.77 25.97 43.09 26.09 0.00 27.00 51.85 50.83 31.33
DuQuant + LWC 4.4e4 26.28 27.10 38.38 25.81 0.00 28.00 50.49 50.12 30.77
BoA 17.31 27.82 47.52 63.27 44.54 27.27 29.40 61.86 54.54 44.53
W3A3KV16 OmniQuant 2.7e4 26.79 26.05 38.62 25.43 0.00 26.20 50.16 50.36 30.45
AffineQuant 3.0e4 26.45 25.25 38.62 25.20 0.00 30.00 50.76 49.41 30.71
SpinQuant-RTN 21.27 27.05 42.93 62.75 45.02 27.07 29.00 62.02 52.25 43.51
SpinQuant-GPTQ 20.49 26.96 40.40 54.77 50.47 26.17 31.80 62.13 53.35 43.26
DuQuant 280.8 23.29 29.42 41.41 30.84 1.37 26.40 51.09 51.22 31.88
DuQuant + LWC 783.1 22.35 27.57 38.07 28.05 0.18 27.00 51.74 50.59 30.69
BoA 17.56 31.66 45.88 57.98 53.55 32.56 32.00 65.34 52.17 46.39
W4A4KV16 OmniQuant 142.4 22.44 35.02 41.65 33.69 2.46 27.60 54.90 51.14 33.61
AffineQuant 144.5 25.43 32.70 45.17 33.72 2.83 25.20 56.58 49.96 33.95
SpinQuant-RTN 8.229 45.14 72.10 75.44 74.07 59.53 40.80 76.44 67.88 63.93
SpinQuant-GPTQ 7.636 46.33 72.31 75.17 72.69 64.27 42.80 76.44 68.27 64.79
DuQuant 7.793 45.90 71.46 72.60 74.36 66.06 42.40 78.02 69.61 65.05
DuQuant + LWC 8.066 44.20 71.13 74.01 73.59 58.55 40.60 76.33 66.77 63.15
BoA 7.496 47.10 72.05 74.98 74.22 66.29 42.00 76.33 69.53 65.31
W4A4KV4 OmniQuant 188.2 22.78 32.32 43.39 31.69 2.01 25.40 54.79 50.20 32.82
AffineQuant 221.5 22.61 31.23 40.98 31.06 1.53 24.60 55.77 50.28 32.26
SpinQuant-RTN 8.503 42.06 69.40 72.78 72.69 60.50 38.40 74.21 65.19 61.90
SpinQuant-GPTQ 7.869 45.56 73.32 71.07 73.88 63.13 39.80 77.04 68.59 64.05
DuQuant 8.000 45.99 70.88 74.01 73.67 64.32 41.00 76.55 68.90 64.42
DuQuant + LWC 8.402 44.80 70.29 73.82 73.69 58.61 39.00 75.08 66.77 62.76
BoA 7.705 45.22 74.54 72.81 74.09 65.72 43.60 77.53 66.69 65.03
  • $\dagger$
    BoA has been applied after transforming the model via SpinQuant.
    The LET option has been deactivated for OmniQuant and AffineQuant because this option does not support models exploiting GQA.
Table 13: Weight-activation quantization performance on the transformed LLaMA3.2-1B
Precision Method Wiki2 PPL ($\downarrow$) Zero-shot Accuracy ($\uparrow$)
Arc-c Arc-e BQ HS LAMB OBQA PIQA WG Average
FP16 Baseline 13.15 38.14 63.26 69.51 60.78 54.38 34.60 74.37 59.51 56.82
W2A4KV16 OmniQuant 2.5e3 24.83 26.85 37.83 26.01 0.00 29.00 52.34 51.78 31.08
AffineQuant 6.1e3 25.00 26.73 37.89 25.64 0.00 26.80 52.18 50.36 30.58
SpinQuant-RTN 93.90 23.63 32.62 54.59 28.71 7.51 24.60 53.43 50.75 34.48
SpinQuant-GPTQ 104.4 21.59 32.28 50.83 30.11 7.55 26.20 53.75 49.96 34.03
DuQuant 1.0e5 26.28 25.04 46.61 26.39 0.00 29.40 49.13 47.51 31.30
DuQuant + LWC 1.0e4 26.88 24.54 38.04 25.78 0.00 28.00 50.98 50.43 30.58
BoA 59.95 23.98 37.04 54.68 32.02 9.33 27.00 56.04 52.41 36.56
W3A3KV16 OmniQuant 7.1e3 26.79 26.60 38.20 25.24 0.00 29.20 51.03 48.93 30.75
AffineQuant 5.6e3 27.39 26.77 38.38 25.59 0.00 29.40 49.89 49.09 30.81
SpinQuant-RTN 57.43 24.23 33.75 53.76 32.59 12.11 25.40 53.86 49.72 35.68
SpinQuant-GPTQ 57.52 23.38 36.07 48.93 35.13 10.89 27.40 55.82 47.12 35.59
DuQuant 2.9e4 25.09 26.14 40.00 25.94 0.01 28.40 50.49 50.75 30.85
DuQuant + LWC 1.2e4 25.17 25.21 39.33 25.65 0.00 27.20 51.85 51.22 30.70
BoA 48.47 23.81 38.97 53.21 35.55 13.90 26.80 54.84 52.49 37.45
W4A4KV16 OmniQuant 156.0 24.23 32.66 47.89 31.15 0.84 27.20 54.84 50.12 33.62
AffineQuant 155.8 23.63 35.14 47.86 30.61 0.97 28.80 54.52 48.30 33.73
SpinQuant-RTN 17.66 32.25 54.80 63.30 53.38 40.69 30.80 68.93 53.75 49.74
SpinQuant-GPTQ 16.68 33.79 55.56 64.77 55.08 41.35 33.40 68.28 54.85 50.89
DuQuant 2.3e4 26.37 26.47 40.28 26.40 0.01 30.00 48.59 48.38 30.81
DuQuant + LWC 1.9e4 26.19 26.81 40.31 25.48 0.00 26.00 50.76 47.12 30.33
BoA 16.25 33.62 58.12 65.72 55.06 44.28 31.80 69.80 55.64 51.76
W4A4KV4 OmniQuant 219.2 23.55 31.82 45.47 29.46 0.79 28.00 51.52 48.22 32.35
AffineQuant 222.0 23.46 32.74 48.53 29.17 0.64 30.20 52.88 52.64 33.78
SpinQuant-RTN 19.68 32.25 52.86 61.71 50.74 34.91 31.60 66.65 53.43 48.02
SpinQuant-GPTQ 18.31 30.72 55.18 61.90 52.91 39.13 31.80 67.30 51.93 48.86
DuQuant 2.0e4 25.60 25.97 39.91 26.03 0.00 27.80 49.02 50.36 30.59
DuQuant + LWC 1.7e4 27.05 26.01 38.96 26.07 0.01 28.40 49.73 49.41 30.71
BoA 17.83 32.94 56.02 64.16 53.16 40.41 31.60 68.93 56.04 50.41
  • $\dagger$
    BoA has been applied after transforming the model via SpinQuant.
    The LET option has been deactivated for OmniQuant and AffineQuant because this option does not support models exploiting GQA.
Table 14: Weight-activation quantization performance on the transformed LLaMA3.2-3B
Precision Method Wiki2 PPL ($\downarrow$) Zero-shot Accuracy ($\uparrow$)
Arc-c Arc-e BQ HS LAMB OBQA PIQA WG Average
FP16 Baseline 11.04 46.16 67.80 78.62 70.44 62.15 36.00 75.52 67.40 63.01
W2A4KV16 OmniQuant 5.8e3 25.94 26.81 37.89 25.83 0.00 28.80 50.44 49.80 30.69
AffineQuant 5.6e3 23.89 26.64 38.41 25.74 0.00 26.00 49.56 49.49 29.97
SpinQuant-RTN 46.31 24.06 31.14 51.87 31.41 12.70 26.60 55.39 50.99 35.52
SpinQuant-GPTQ 68.74 23.12 32.45 38.29 31.84 9.53 27.20 54.95 51.30 33.59
DuQuant 1.2e5 25.94 24.75 39.91 26.77 0.00 29.20 51.63 49.80 31.00
DuQuant + LWC 1.2e3 24.91 25.63 37.83 26.16 0.00 28.40 52.72 48.07 30.47
BoA 34.25 24.57 36.28 60.24 38.57 17.38 29.40 57.56 52.49 39.56
W3A3KV16 OmniQuant 1.2e4 25.00 26.56 37.95 25.84 0.00 27.20 49.73 51.14 30.43
AffineQuant 1.0e4 25.77 25.72 38.13 26.21 0.00 28.20 50.05 49.72 30.48
SpinQuant-RTN 36.37 24.23 35.52 55.66 36.04 15.19 26.80 57.78 49.72 37.62
SpinQuant-GPTQ 27.23 28.33 40.70 60.73 44.15 21.53 28.20 58.16 54.06 41.98
DuQuant 564.6 22.01 27.02 43.21 29.68 0.94 27.00 50.60 52.80 31.66
DuQuant + LWC 82.65 23.04 31.61 48.75 33.24 3.01 28.80 53.32 53.43 34.40
BoA 23.76 28.75 44.53 60.98 46.08 29.32 28.40 60.07 54.22 44.04
W4A4KV16 OmniQuant 131.8 24.57 34.55 48.75 36.10 3.21 27.00 56.37 51.07 35.20
AffineQuant 131.8 23.55 34.60 47.65 35.80 3.01 27.80 56.26 49.57 34.78
SpinQuant-RTN 12.42 38.23 60.94 72.69 64.68 55.02 31.60 71.22 61.09 56.93
SpinQuant-GPTQ 11.87 39.51 63.05 74.31 66.42 56.57 35.80 71.22 62.83 58.71
DuQuant 13.91 38.40 59.93 74.59 66.43 53.84 36.60 70.73 60.46 57.62
DuQuant + LWC 13.32 38.23 60.73 75.29 65.93 51.58 36.40 71.87 63.38 57.93
BoA 11.57 40.70 63.68 75.50 66.86 56.77 36.00 70.24 63.61 59.17
W4A4KV4 OmniQuant 169.1 23.89 34.22 43.76 32.36 1.62 26.20 54.24 50.51 33.35
AffineQuant 197.1 21.42 32.87 42.20 31.70 1.25 27.00 55.22 50.75 32.80
SpinQuant-RTN 12.96 39.16 61.87 72.42 63.06 48.84 34.40 69.70 58.80 56.03
SpinQuant-GPTQ 12.24 39.33 61.91 72.32 65.56 54.57 35.20 70.13 61.33 57.54
DuQuant 14.65 39.51 60.02 72.08 65.76 52.54 34.00 69.53 62.19 56.95
DuQuant + LWC 13.84 38.99 59.68 72.05 64.76 49.18 35.60 70.40 61.56 56.53
BoA 11.98 39.08 63.64 74.22 65.60 55.51 38.80 71.22 63.14 58.90
  • $\dagger$
    BoA has been applied after transforming the model via SpinQuant.
    The LET option has been deactivated for OmniQuant and AffineQuant because this option does not support models exploiting GQA.
Appendix D Pseudocode for GPTQ
In this appendix, we provide the pseudocode of the conventional GPTQ (Frantar et al., 2023), which is omitted in the main manuscript due to the page limitation.
Algorithm 2 GPTQ
0:  weights $\mathbf{W}$, Hessian information $\mathbf{U}_{\text{col}}$, and pre-determined step size $\mathbf{S}$
1:  Initialize quantized output: $\mathbf{Q}\leftarrow\mathbf{0}_{d_{\text{row}}\times d_{\text{col}}}$
2:  Initialize quantization errors: $\mathbf{E}\leftarrow\mathbf{0}_{d_{\text{row}}\times d_{\text{col}}}$
3:  for $j=0,\cdots,d_{\text{col}}-1$ do
4:     Quantize the $j$-th column: $\mathbf{Q}_{:,j}\leftarrow\text{quant}(\mathbf{W}_{:,j},\mathbf{S})$
5:     Estimate quantization error: $\mathbf{E}_{:,j}\leftarrow(\mathbf{W}_{:,j}-\mathbf{Q}_{:,j})/[\mathbf{U}_{\text{col}}]_{j,j}$
6:     Update weights: $\mathbf{W}_{:,j:}\leftarrow\mathbf{W}_{:,j:}-\mathbf{E}_{:,j}\cdot[\mathbf{U}_{\text{col}}]_{j,j:}$
7:  end for
7:  quantized weights $\mathbf{Q}$, quantization error $\mathbf{E}$